### v1

# Behavior of prediction performance metrics with rare events

Abstract
Area under the receiving operator characteristic curve (AUC) is commonly reported alongside binary prediction models. However, there are concerns that AUC might be a misleading measure of prediction performance in the rare event setting. This setting is common since many events of clinical importance are rare events. We conducted a simulation study to determine when or whether AUC is unstable in the rare event setting. Specifically, we aimed to determine whether the bias and variance of AUC are driven by the number of events or the event rate. We also investigated the behavior of other commonly used measures of prediction performance, including positive predictive value, accuracy, sensitivity, and specificity. Our results indicate that poor AUC behavior—as measured by empirical bias, variability of cross-validated AUC estimates, and empirical coverage of confidence intervals—is driven by the minimum class size, not event rate. Performance of sensitivity is driven by the number of events, while that of specificity is driven by the number of non-events. Other measures, including positive predictive value and accuracy, depend on the event rate even in large samples. AUC is reliable in the rare event setting provided that the total number of events is moderately large.
Funding
The authors disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: Research reported in this publication was supported by the National Institute of Mental Health of the National Institutes of Health under award numbers R01 MH125821, U19-MH099201, and U19-MH121738.
Acknowledgments
The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Declaration of competing interest
The authors declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Ethical approval and informed consent
No new data were created for this study. The data analyzed in this study were obtained under approval from the appropriate institutional review boards for the participating health systems.
Data availability
The datasets generated and analyzed during this study are not publicly available because they contain detailed information from the electronic health records in the health systems participating in this study and are governed by the Health Insurance Portability and Accountability Act (HIPAA). Data are however available from the authors upon reasonable request, with permission of all health systems involved and fully executed data use agreement.
1 Introduction
Clinical prediction models can be used to guide clinical decision-making through early warning systems (Patton and Liu, 2023), identify patients at risk for an adverse outcome (Simon et al., 2018), obtain accurate prognoses and diagnoses (Obermeyer and Emanuel, 2016), and better allocate health resources (Chen and Asch, 2017). When the outcome of interest is binary, prediction performance can be measured in a variety of ways, including sensitivity and specificity (see, e.g., Metz, 1978), Brier score (see, e.g., Steyerberg et al., 2001), and the area under the receiver operating characteristic curve (AUC; see, e.g., Hanley and McNeil, 1982). Often, assessing several prediction performance metrics can provide a full picture of model performance, but focus on any one metric depends on how the model will ultimately be used (Cook, 2007; Pepe et al., 2007; Janket et al., 2007).
Estimating and reporting the performance of risk prediction models in the rare event setting presents several challenges. First, even with a large number of observations, the absolute number of events may be small. In predictive modeling of binary outcomes, the number of events, not just the total sample size, is of importance (Vergouwe et al., 2005). Additionally, a small event rate impacts the interpretation of some common measures of predictive model performance because the overall event rate impacts the behaviors of these measures (Lever et al., 2016; Saito and Rehmsmeier, 2015). There has been some recent concern that AUC, and receiver operating characteristic curves more generally, are misleading or uninformative in the rare event setting (Saito and Rehmsmeier, 2015; Lever et al., 2016; Steyerberg et al., 2018; Adhikari et al., 2021; Bokhari, 2023). In this manuscript, we aimed to determine whether a low event rate regardless of sample size or a small number of events alone leads to either instability in AUC estimates or AUC estimates that do not reflect the true AUC.
Our motivation for considering AUC in the rare-event setting is that the AUC, among other prediction performance metrics, is often used in the literature to describe the performance of suicide risk prediction models. In our collaborations developing suicide risk prediction models (Simon et al., 2018; Coley et al., 2021; Shortreed et al., 2023), the per-visit 90-day suicide attempt rate is close to 0.65% for visits to mental health specialty clinics and 0.33% for general medical visits with mental health diagnoses; suicide deaths are even more rare, approximately 0.02% and 0.01% for mental health specialty visits and general medical visits, respectively. Despite this rarity, suicide was the 10th leading cause of death in the United States in 2019; there were over suicide deaths, representing suicide deaths per population (Xu et al., 2021). Predicting suicide deaths and attempts in order to identify individuals at risk—so that these individuals can be offered interventions—is clearly of clinical importance. A 2019 systematic review found a total of 64 unique suicide prediction models (Belsher et al., 2019). While suicide attempt is rare, the total sample size for developing these prediction models can be large. Thus, having a clear understanding of the behavior of AUC and other prediction metrics in the rare-event setting is important.
While the focus of the experiments presented here is AUC, we also investigated the behavior of multiple other metrics: sensitivity, specificity, positive predictive value (PPV) and negative predictive value (NPV), the F score (Van Rijsbergen, 1974), classification accuracy, and the Brier score. Each of these metrics depends differently on event rate, the number of events, and the number of non-events, and thus, may behave differently than AUC.
The remainder of the article is organized as follows. In the next section, we provide explicit definitions of the prediction performance metrics we consider. In Section 3, we describe our simulation methods, and present the results in Section 4. We provide concluding remarks in Section 5. Results of additional simulations, and results for several performance metrics, can be found in the Supplementary Material.
2 Prediction performance metrics: notation and definitions
We suppose that we have independent and identically distributed observations drawn from a distribution , where the data unit is ; is a covariate vector and is the binary outcome of interest. Suppose further that we have a prediction model that has been trained on an independent sample drawn from (e.g., a training sample or a set of folds in a cross-validation procedure) and that returns a predicted probability of the outcome based on the observed covariates . Then, the AUC of prediction model with respect to distribution is
where and are independent draws from . In other words, the AUC is the probability that the predicted probability is correctly ordered lower for a randomly sampled non-event than for a randomly sampled event. The AUC can take values between 0.5 and 1; an AUC of 0.5 implies that the prediction model performs no better than assigning predicted probabilities at random (Bradley, 1997), while an AUC of one implies that the predicted probability of an event is always larger than the predicted probability of a non-event.
Next, we define sensitivity, specificity, PPV and NPV, the F score (Van Rijsbergen, 1974), classification accuracy, and the Brier score. All metrics besides the Brier score are defined relative to a threshold, , of predicted risk. The sensitivity (sens) and specificity (spec) of given threshold with respect to are
that is, sensitivity is the probability that the model-predicted probability of an event is higher than threshold given that the observation truly is an event and specificity is the probability that the model-predicted probability of an event is lower than threshold given that the observation truly is a non-event.
The positive and negative predictive values of given threshold relative to are
in other words, PPV is the probability that an observation is an event given the model-predicted probability of an event is greater than threshold and NPV is the probability that an observation is a non-event given the model-predicted probability of an event is less than threshold .
The F score of model given threshold with respect to trades off sensitivity and PPV; where the amount of the trade off is dictated by a constant, :
Two common choices are and , resulting in
The score weights sensitivity and PPV equally, while the score places higher weight on PPV; this can be useful in settings where there is a high cost to intervening based on . In many rare outcome settings the potential interventions under consideration are relatively high cost, thus putting a larger weight on PPV assures that we will be intervening on individuals at higher risk, potentially at a cost of missing some events.
The classification accuracy of given threshold with respect to is
where is the indicator function, i.e., if is true and otherwise. In other words, accuracy is the probability that the model-predicted event status defined using threshold is equal to the true event status, giving equal weight to events and non-events (Jiang et al., 2021).
Finally, the Brier score of with respect to is
where is shorthand for the expectation under . In other words, the Brier score is the mean squared error of .
3 Simulation methods
We now outline the methods used to investigate the behavior of each these performance metrics in rare-event settings, with both large and small sample size.
3.1 Data
There has been growing interest in estimating risk prediction models for self-harm and suicide, important and rare health outcomes. This work evaluating the behavior of prediction model performance metrics is motivated by our collaborations estimating suicide risk prediction models using health records data (see, e.g., Simon et al., 2018; Coley et al., 2021). To maintain both the complex outcome-predictor relationships and the complex correlation structure seen between predictors in our motivating dataset, we used a plasmode simulation approach (see, e.g., Schreck et al., 2024). Specifically, we constructed a “parent” dataset with independent observations by randomly sampling one mental health visit per patient from 25 million outpatient mental health visits across seven Mental Health Research Network sites (https://mhresearchnetwork.org). These visits were made by patients aged 13 years or older and occurred between January 1, 2009 and September 30, 2017 (Coley et al., 2021). For patients with an event at any visit—defined as non-fatal or fatal self-inflicted injury or poisoning within 90 days of a mental health visit—a visit with an event was sampled. The event rate in the resulting dataset, , was 0.92%. There were 149 predictors, consisting of demographic characteristics, mental health and substance abuse diagnoses (including past self-inflicted injuries or poisonings), other medical diagnoses, past inpatient or emergency mental health care, dispensed mental health medications, the patient health questionnaire (PHQ) 8-item score (PHQ-8) representing depressive symptoms, and the PHQ 9th item response indicating thoughts of death or self-harm (Coley et al., 2021; Simon et al., 2018). All predictors except patient age in years and PHQ-8 score (0–24) were coded as binary indicators; the PHQ item 9, which has 4 possible responses (0,1,2,3), was coded using 3 categorical variables indicating each possible non-zero response.
3.2 Simulation scenarios
In our simulations, we bootstrap sampled the covariates with replacement from the parent dataset with observations. Then we generated outcomes according to a fitted regression model estimated in that original dataset (see, e.g., Franklin et al., 2017) with varying event rates. To do this, we sampled three datasets of size = 1 million with event rates %, %, and %. In each of these data sets, we fit logistic regression, ridge regression (Hoerl and Kennard, 1970), and random forest models (Breiman, 2001). These became three outcome-generating functions for our simulations for the three event rates. The tuning parameters used to fit each outcome model are provided in Table S1; 10-fold cross-validated estimates of the performance metric values on the appropriate dataset are presented in Table 1.
| Algorithm | Event rate | AUC | Brier | Percentile | Sens. | Spec. | Acc. | PPV | NPV | F1 | F0.5 |
| 90 | 0.613 | 0.876 | 0.875 | 0.022 | 0.998 | 0.043 | 0.028 | ||||
| 95 | 0.503 | 0.93 | 0.928 | 0.032 | 0.998 | 0.06 | 0.039 | ||||
| 0.005 | 0.806 | 0.004 | 99 | 0.263 | 0.989 | 0.985 | 0.097 | 0.997 | 0.141 | 0.111 | |
| 90 | 0.65 | 0.872 | 0.87 | 0.045 | 0.996 | 0.084 | 0.055 | ||||
| 95 | 0.534 | 0.936 | 0.932 | 0.072 | 0.995 | 0.127 | 0.087 | ||||
| 0.009 | 0.832 | 0.008 | 99 | 0.194 | 0.996 | 0.988 | 0.296 | 0.993 | 0.234 | 0.268 | |
| 90 | 0.671 | 0.881 | 0.877 | 0.095 | 0.993 | 0.167 | 0.115 | ||||
| 95 | 0.536 | 0.945 | 0.937 | 0.154 | 0.991 | 0.239 | 0.18 | ||||
| RF | 0.018 | 0.854 | 0.015 | 99 | 0.144 | 0.998 | 0.982 | 0.568 | 0.984 | 0.229 | 0.357 |
| 90 | 0.642 | 0.902 | 0.901 | 0.029 | 0.998 | 0.056 | 0.036 | ||||
| 95 | 0.517 | 0.952 | 0.95 | 0.047 | 0.998 | 0.087 | 0.058 | ||||
| 0.005 | 0.865 | 0.004 | 99 | 0.257 | 0.991 | 0.988 | 0.118 | 0.997 | 0.162 | 0.132 | |
| 90 | 0.65 | 0.872 | 0.87 | 0.045 | 0.996 | 0.084 | 0.055 | ||||
| 95 | 0.534 | 0.936 | 0.932 | 0.072 | 0.995 | 0.127 | 0.087 | ||||
| 0.009 | 0.008 | 99 | 0.194 | 0.996 | 0.988 | 0.296 | 0.993 | 0.234 | 0.268 | ||
| 90 | 0.632 | 0.91 | 0.905 | 0.116 | 0.993 | 0.196 | 0.139 | ||||
| 95 | 0.494 | 0.958 | 0.95 | 0.181 | 0.99 | 0.265 | 0.208 | ||||
| Ridge | 0.018 | 0.866 | 0.016 | 99 | 0.219 | 0.994 | 0.98 | 0.404 | 0.986 | 0.284 | 0.346 |
| 90 | 0.647 | 0.902 | 0.901 | 0.03 | 0.998 | 0.057 | 0.037 | ||||
| 95 | 0.517 | 0.952 | 0.95 | 0.047 | 0.998 | 0.087 | 0.058 | ||||
| 0.005 | 0.867 | 0.004 | 99 | 0.257 | 0.991 | 0.988 | 0.117 | 0.997 | 0.161 | 0.132 | |
| 90 | 0.643 | 0.905 | 0.903 | 0.059 | 0.996 | 0.108 | 0.072 | ||||
| 95 | 0.509 | 0.954 | 0.95 | 0.093 | 0.995 | 0.158 | 0.112 | ||||
| 0.009 | 0.008 | 99 | 0.247 | 0.992 | 0.985 | 0.228 | 0.993 | 0.237 | 0.232 | ||
| 90 | 0.635 | 0.91 | 0.905 | 0.117 | 0.993 | 0.197 | 0.139 | ||||
| 95 | 0.494 | 0.958 | 0.95 | 0.181 | 0.99 | 0.265 | 0.207 | ||||
| GLM | 0.018 | 0.868 | 0.016 | 99 | 0.22 | 0.994 | 0.98 | 0.405 | 0.986 | 0.286 | 0.347 |
Abbreviations: Sens: sensitivity; Spec: specificity; Acc: accuracy; GLM: logistic regression; RF: random forests; Ridge: ridge regression.
To investigate the behavior of prediction performance metrics, we varied both the event rate (keeping sample size fixed) and the absolute number of events (increasing sample size appropriately as the event rate decreases). We varied and considered the three event rates defined above. For each scenario, we first bootstrap sampled with replacement covariates for observations (for a training dataset) and, separately, covariates for 1.6 million observations (for an evaluation dataset). The evaluation dataset is a large sample of completely independent data; performance in this dataset is intended to capture the true model performance in the population. Once we sampled covariates for each of the datasets (training, evaluation), using each of the three algorithms we generated outcomes from the appropriate model to obtain each event rate. For each scenario defined by event rate, sample size, and algorithm, we generated 2500 random datasets following the procedure described above (i.e., we conducted 2500 simulation iterations).
3.3 Developing clinical prediction models
Estimation of prediction models within each training set used the same three prediction algorithms described above for generating outcomes: logistic regression, ridge logistic regression, and random forests. While logistic regression is a common prediction approach in binary outcome settings, in some rare-event cases, it may be unstable, particularly with a large number of predictors. Ridge regression was selected as an example of a penalized regression method that might be used in a setting with a large number of predictors (Hoerl and Kennard, 1970). Random forests were selected as an example of a more complex prediction modeling approach that allow for interactions between predictors without being explicitly coded, in contrast to logistic or ridge regression. We used probability trees (Malley et al., 2012) within random forests. In our simulations, we did not include any interactions between our predictors for logistic regression or ridge regression.
Ridge regression and random forest require the specification of several key tuning parameters: the regularization parameter for ridge, and the minimum node size, number of predictors, , considered at each split, the depth of tree, and number of trees for random forests. For random forests, we fixed (the square root of the number of predictors) and allowed tree depth to be controlled by minimum node size. We selected global tuning parameters that were used in all simulation replicates in a small simulation study with 100 replicates. In this small simulation study, for each simulation replication, we used 10-fold cross-validation to obtain optimal tuning parameters. Details on this simulation are provided in Section S1.5 of the Supplementary Materials. The final tuning parameters used for each algorithm in the main simulations presented here are provided in Table S2. We used global tuning parameters, because the goal of our simulation study was not to optimize the prediction performance of a given model but rather to understand how the behavior of prediction model performance measures depended on event rate and absolute number of events. We did not need the tuning parameters to be the optimal values for each sampled replication-specific training set, since we were interested in the performance of a model estimated with particular values of tuning parameters.
3.4 Assessing clinical prediction model performance
To provide an accurate estimate of prediction performance based solely on the training data, we used cross-validation. The training set was divided into ten folds stratified by the outcome, and 10-fold cross-validation was performed to estimate AUC, sensitivity, specificity, PPV, NPV, accuracy, , , and Brier score. At the smaller training set sizes and event rates, on occasion folds had no events despite using outcome-stratified sampling (i.e., there were fewer than 10 events in the sampled training data). These folds without events were excluded during cross-validation, since at least one true event was required to obtain an estimate of the hold-out fold AUC, and AUC was the primary prediction performance metric of interest. The metrics requiring thresholds (all except AUC and Brier score) were calculated at each of the , , and percentiles of predicted probabilities from the training folds.
We used empirical standard errors (ESEs) to describe the variability of performance measure metrics. Code implementing asymptotic standard errors (ASEs) were not available for many metrics. While the ESE is not available in a data analysis, in the simulation setting it provides a best-case benchmark for variance estimation if we were able to consistently estimate the standard error for each performance metric. For each performance metrics the ESE was defined as the standard deviation of the point estimates across the 2500 simulation replications. For AUC, in addition to the ESE, we calculated influence curve based confidence intervals (asymptotic standard errors, or ASEs) (LeDell et al., 2015), as code was readily available.
We used the evaluation dataset to estimate the true prediction performance of each of the three algorithms in our population of interest. We fit each of the three algorithms on the entire training set in each simulation replication to obtain a fitted model and corresponding thresholds of predicted probabilities , , and . We computed all prediction performance metrics on the evaluation dataset using this fitted model and thresholds for each replication. The final “true" value of each prediction performance metric was then defined as the average of the evaluation-set values across the 2500 simulation replications.
For each performance metric and each simulation scenario, we measured reliability by calculating the empirical bias, variance, and mean squared error (MSE) of the estimated performance metric, using the evaluation-estimated mean performance metric as the gold standard. We calculated empirical coverage of the 95% confidence intervals based on the ASE (for AUC) and on the ESE (for all metrics), comparing the confidence interval from each replication with the evaluation-estimated mean performance metric. The computation of these reliability measures is provided in Table 2.
| Performance measure | Formula or measure(s) |
| Empirical bias | |
| Empirical standard error (ESE) | |
| Empirical coverage of 95% CI | |
| Empirical mean squared error |
3.5 Computing
All simulations and analyses were conducted in R version 4.1.1 (R Core Team, 2013). Logistic regression was implemented using the base R stats package, penalized regression using the package glmnet (version 4.1-4) (Friedman et al., 2010), and random forests using the package ranger (version 0.14.1) (Wright and Ziegler, 2017). Influence curve-based confidence intervals for cross-validated AUC were obtained using the cvAUC package (LeDell et al., 2022). The foreach package was used to implement parallel computing (Daniel et al., 2022). For reproducibility, each replication of the simulations used a unique random number generation seed from a pre-defined sequence of random seeds.
4 Simulation results
We define the effective sample size for a prediction metric as the amount of information for estimating that metric in the entire training sample. For sensitivity, the effective sample size is the number of events in the training sample; for specificity, it is the number of non-events in the training sample. For AUC, the effective sample size is the minimum of the number of events and non-events; because the event rate in our scenario is less than 50%, the effective sample size is the number of events.
The true value of each performance measure depends on the sample size, event rate, and model-estimation algorithm. In Table 3, we show the true values of AUC in all settings; we present the true values for other estimands in Table LABEL:tab:all_true_values_test.
| Event rate = 0.009 | Event rate = 0.018 | Event rate = 0.005 | |||||||
| n | GLM | RF | Ridge | GLM | RF | Ridge | GLM | RF | Ridge |
| 0.571 | 0.8 | 0.813 | 0.656 | 0.799 | 0.807 | 0.533 | 0.768 | 0.817 | |
| 0.669 | 0.806 | 0.819 | 0.745 | 0.8 | 0.814 | 0.565 | 0.796 | 0.823 | |
| 0.83 | 0.821 | 0.831 | 0.835 | 0.816 | 0.824 | 0.809 | 0.813 | 0.836 | |
| 0.846 | 0.826 | 0.834 | 0.842 | 0.82 | 0.827 | 0.843 | 0.817 | 0.841 | |
| 0.855 | 0.833 | 0.839 | 0.847 | 0.824 | 0.83 | 0.863 | 0.832 | 0.847 | |
Abbreviations: GLM: logistic regression; RF: random forests; Ridge: ridge regression.
We display results describing the behavior of AUC estimates in Figure 1. Bias (top panel) for all algorithms decreases to zero as the effective sample size increases. The effective sample size for the AUC is driven by the number of events, i.e., for event rate , such that the three right farthest dots in each panel show the AUC bias at each event rate at a sample size of 1,000,000. We can see that within each sample size (shown with difference colors in Figure 1) as the event rate increases the bias decreases, but that once a minimum number of events is reached (approximately 1837 events), the sample size or event rate no longer impacts the bias, as bias is effectively zero. In Figure S1, we show bias and coverage versus training set size for each event rate; the trends in bias and coverage do not vary with event rate. Taken together, this shows that it is the effective sample size, not the event rate, that is driver of AUC bias. This is supported by the form of the influence function for AUC derived by LeDell et al. (2015, Theorem 4.1): the estimated AUC (and its asymptotic variance) depends on the number of events and number of non-events, and will be impacted if one of these is small.
Coverage of 95% confidence intervals (bottom panel of Figure 1) based on the ESE is near the nominal level at all sample sizes (denoted with dots in Figure 1), while coverage of ASE-based intervals followed the same pattern as bias, increasing to the nominal level for all event rates and algorithms by an effective sample size of 1837. In Figure S2, we show mean squared error and confidence interval width for AUC, showing that the variance decreases with increasing effective sample size. We expect poor coverage of influence curve-based (ASE) intervals when the effective sample size is small, since these intervals use an asymptotic estimate of the variance of the cross-validated AUC and thus are asymptotic 95% intervals (LeDell et al., 2015). We observed poor behavior (bias and low coverage) for estimating AUC using logistic regression when the number of predictors exceeded (or was similar to) the effective sample size; this is consistent with work that has shown increased bias and variability in parameter estimates when the number of events per variable is low. A commonly suggested guideline, and one supported by these simulation studies, is that there should be ten or more events per variable to estimate a logistic regression model (Peduzzi et al., 1996).
The performance of sensitivity and specificity are also driven by the effective sample size. Figure 2 shows sensitivity and specificity using the 95th percentile of predicted risk, while results for sensitivity and specificity at the 90th and 99th percentile following a similar pattern are found in Tables S4–S7 in the Supplementary Materials. As with AUC, the bias decreases with increasing effective sample size; coverage based on the ESE is at the nominal level for sensitivity. For specificity, while bias follows the same pattern, coverage follows a slightly different pattern. While coverage is mostly near the nominal level, when the true specificity is close to 1 (as is the case for specificity in general but particularly at the 99th percentile), there is instability of interval coverage estimates.
In contrast to AUC and sensitivity, the performance of accuracy, PPV, NPV, , , and Brier score relied in part on the event rate itself, rather than just on the effective sample size. We display the results for PPV at the 95th percentile of predicted risk in Figure 3; performance for accuracy, NPV, , , and Brier score and the other algorithms is similar to the results for PPV and is provided in the Supplementary Material (Tables S8–S22, which also include all metrics at the 90th and 99th percentiles). For PPV, at a given training set size, the bias is larger for smaller event rates; this is most noticeable at the smallest training set sizes (5,000 and 10,000). Coverage was again at the nominal level. Similar results hold for the other estimands: for a given sample size, bias and coverage tended to be worse at the lower event rates (Tables S8–S22).
5 Discussion
We found through numerical experiments that the behavior (bias and confidence interval coverage) of AUC, sensitivity, and specificity is driven by the effective sample size, not the event rate. Specifically, once a minimum number of events was reached through either a larger sample size or a higher event rate, estimates of AUC and sensitivity estimates stabilized. For specificity, the effective sample size is the number of non-events and thus this stability was often reached early as we investigated rare event settings. In our simulations, as the effective sample size grew, regardless of the event rate, the cross-validated AUC converged to true AUC and empirical coverage of confidence intervals based on the influence function or the empirical standard error was close to the nominal level. Even in the setting of a very rare event—the smallest event rate we considered was , or approximately one event for every 217 non-events—these simulations show that AUC performs well (in terms of low bias for estimating the true AUC) provided that the number of events in the training set is sufficiently large. The largest training set size we considered was 1 million, which corresponds to 4,591 training set events for the smallest event rate.
While the effective sample size is a function of the event rate, it is also a function of the training set size. This implies that in large sample settings, even with a very rare event we will still have sufficiently many events for the estimated prediction metric to provide a reliable understanding of the performance of a predictive model. We do not have to worry about these prediction metrics (AUC, sensitivity, and specificity) being unreliable simply because the event rate is small. While this distinction between poor behavior of a prediction metric with respect to its reliability in rare event settings may be unimportant in settings where large data sets are not available, it is crucial in settings—such as the use of health-system electronic health records data to build clinical prediction models—where large data sets are increasingly available and computationally feasible to work with.
In contrast to AUC, the behavior of accuracy, Brier score, PPV, NPV, score, and score all showed some direct reliance on event rate in addition to being driven by the size of the training set. This is not surprising, since the marginal outcome probability is included in the definition of each of these metrics. However, bias converged to zero and coverage approached the nominal level as the effective sample size grew, regardless of event rate, indicating that the behavior of these metrics relies on both the effective sample size and the training data size.
The reliance of performance metrics on effective sample size emphasizes the need to choose an appropriate prediction modeling approach for the data. This may include dimension reduction in some cases. For example, in a situation with an effective sample size of 23 (the smallest we considered in our simulations), we would not expect a logistic regression model with 150 predictors to perform well without some dimension reduction; indeed, we observed poor performance for logistic regression in this scenario.
There are two main limitations to our work. The first is that we did not vary the strength of association between outcome and predictors in our simulations: it is possible that results would be improved at a smaller effective sample size if there are strong predictors of the outcome. However, our results still provide useful information about the performance of these metrics in a scenario with complex relationships between variables. The second is that we do not have asymptotic standard errors for metrics besides AUC. While a bootstrap might be tempting in this scenario, the poor performance of empirical standard error-based intervals in some cases, due to the performance metrics being near the boundary values of 0 or 1, suggests that the standard bootstrap may not perform well in these scenarios. Bootstrap procedures designed for constructing intervals when estimates are near the boundary are available and would likely need to be implemented for these measures.
Regardless of the event rate or effective sample size, our results highlight the importance of reporting multiple prediction performance metrics and confidence intervals for these metrics, particularly when the outcome of interest is binary. In the binary outcome setting, different performance metrics (e.g., AUC, sensitivity, PPV) can provide complementary information about model behavior. Others have advocated this approach (see, e.g., Pepe et al., 2007; Hand, 2012; Lever et al., 2016; Adhikari et al., 2021). By reporting multiple metrics and describing their variability, the results can be placed in the proper context.
However, when comparing performance across prediction models and for model selection, it may be necessary to choose a single metric. This choice should be driven by careful consideration of the specific scientific interest and, when applicable, the intended use of the model. For example, if a prediction model is intended to identify patients with a 90-day risk of suicide attempt exceeding 5% (i.e., above the 95th percentile), then an important metric could be sensitivity at the 95th percentile of predicted risk. This work provides us with information on how stable and reliable we can expect our performance metric estimates to be given the event rate and sample size we have on hand and it reassures us that the effective sample size, not the event rate itself, is the factor that limits the reliability of the AUC measure.
SUPPLEMENTARY MATERIALS
S1 Additional simulation results
S1.1 Tuning parameters
In Table S1, we provide the tuning parameters used in the original outcome regression fits used to generate outcomes in the plasmode simulations presented in the main manuscript.
| Event rate | Algorithm | Tuning parameter(s) |
| GLM | — | |
| GLM | — | |
| GLM | — | |
| Ridge | ||
| Ridge | ||
| Ridge | ||
| RF | mtry = 12, ntree = 50, min.node.size = 10,000 | |
| RF | mtry = 12, ntree = 50, min.node.size = 10,000 | |
| RF | mtry = 12, ntree = 50, min.node.size = 10,000 |
In Table S2, we present the tuning parameters used to fit each algorithm at each sample size and event rate in the simulations.
| Training set size | ||||||
| Algorithm | Event rate | |||||
| Ridge | 0.368710 | 0.112562 | 0.007737 | 0.004101 | 0.000887 | |
| Ridge | 0.271674 | 0.056800 | 0.007044 | 0.001855 | 0.001650 | |
| Ridge | 0.137977 | 0.049776 | 0.003318 | 0.002995 | 0.002987 | |
| Random forests | 1000 | 5000 | 5000 | 5000 | 15000 | |
| Random forests | 1000 | 5000 | 5000 | 5000 | 15000 | |
| Random forests | 1000 | 5000 | 5000 | 5000 | 50000 | |
S1.2 True values of estimands
In Table LABEL:tab:all_true_values_test, we present the true values for each estimand at each combination of sample size, event rate, and algorithm.
Abbreviations: GLM: logistic regression; RF: random forests; Ridge: ridge regression.
| Event rate = 0.009 | Event rate = 0.018 | Event rate = 0.005 | ||||||||
| n | Estimand | GLM | RF | Ridge | GLM | RF | Ridge | GLM | RF | Ridge |
| AUC | 0.571 | 0.8 | 0.813 | 0.656 | 0.799 | 0.807 | 0.533 | 0.768 | 0.817 | |
| AUC | 0.669 | 0.806 | 0.819 | 0.745 | 0.8 | 0.814 | 0.565 | 0.796 | 0.823 | |
| AUC | 0.83 | 0.821 | 0.831 | 0.835 | 0.816 | 0.824 | 0.809 | 0.813 | 0.836 | |
| AUC | 0.846 | 0.826 | 0.834 | 0.842 | 0.82 | 0.827 | 0.843 | 0.817 | 0.841 | |
| AUC | 0.855 | 0.833 | 0.839 | 0.847 | 0.824 | 0.83 | 0.863 | 0.832 | 0.847 | |
| Sensitivity 95% | 0.215 | 0.325 | 0.429 | 0.272 | 0.314 | 0.401 | 0.329 | 0.301 | 0.442 | |
| Sensitivity 95% | 0.309 | 0.336 | 0.438 | 0.35 | 0.319 | 0.411 | 0.229 | 0.325 | 0.453 | |
| Sensitivity 95% | 0.458 | 0.354 | 0.46 | 0.44 | 0.332 | 0.429 | 0.454 | 0.353 | 0.479 | |
| Sensitivity 95% | 0.479 | 0.359 | 0.466 | 0.451 | 0.336 | 0.434 | 0.492 | 0.358 | 0.487 | |
| Sensitivity 95% | 0.495 | 0.369 | 0.475 | 0.46 | 0.34 | 0.44 | 0.522 | 0.372 | 0.5 | |
| Specificity 95% | 0.951 | 0.952 | 0.953 | 0.954 | 0.954 | 0.956 | 0.758 | 0.947 | 0.951 | |
| Specificity 95% | 0.953 | 0.953 | 0.953 | 0.956 | 0.955 | 0.956 | 0.952 | 0.951 | 0.952 | |
| Specificity 95% | 0.954 | 0.953 | 0.954 | 0.957 | 0.955 | 0.957 | 0.952 | 0.951 | 0.952 | |
| Specificity 95% | 0.954 | 0.953 | 0.954 | 0.957 | 0.955 | 0.957 | 0.952 | 0.951 | 0.952 | |
| Specificity 95% | 0.954 | 0.953 | 0.954 | 0.957 | 0.955 | 0.957 | 0.952 | 0.952 | 0.952 | |
| PPV 95% | 0.039 | 0.059 | 0.078 | 0.093 | 0.11 | 0.136 | 0.013 | 0.027 | 0.043 | |
| PPV 95% | 0.057 | 0.061 | 0.08 | 0.12 | 0.112 | 0.14 | 0.023 | 0.03 | 0.044 | |
| PPV 95% | 0.084 | 0.065 | 0.084 | 0.15 | 0.117 | 0.146 | 0.044 | 0.033 | 0.046 | |
| PPV 95% | 0.088 | 0.066 | 0.085 | 0.153 | 0.118 | 0.148 | 0.048 | 0.034 | 0.047 | |
| PPV 95% | 0.091 | 0.068 | 0.087 | 0.156 | 0.119 | 0.15 | 0.051 | 0.035 | 0.048 | |
| Sensitivity 90% | 0.29 | 0.47 | 0.539 | 0.368 | 0.458 | 0.516 | 0.409 | 0.438 | 0.552 | |
| Sensitivity 90% | 0.405 | 0.481 | 0.548 | 0.463 | 0.461 | 0.527 | 0.305 | 0.469 | 0.561 | |
| Sensitivity 90% | 0.575 | 0.502 | 0.572 | 0.564 | 0.478 | 0.547 | 0.565 | 0.5 | 0.587 | |
| Sensitivity 90% | 0.596 | 0.511 | 0.578 | 0.575 | 0.489 | 0.553 | 0.605 | 0.507 | 0.596 | |
| Sensitivity 90% | 0.613 | 0.524 | 0.587 | 0.584 | 0.496 | 0.558 | 0.636 | 0.525 | 0.608 | |
| Sensitivity 99% | 0.113 | 0.128 | 0.221 | 0.132 | 0.118 | 0.189 | 0.103 | 0.107 | 0.234 | |
| Sensitivity 99% | 0.156 | 0.14 | 0.231 | 0.165 | 0.126 | 0.198 | 0.117 | 0.129 | 0.249 | |
| Sensitivity 99% | 0.242 | 0.147 | 0.249 | 0.213 | 0.13 | 0.211 | 0.248 | 0.148 | 0.271 | |
| Sensitivity 99% | 0.257 | 0.151 | 0.254 | 0.221 | 0.133 | 0.215 | 0.276 | 0.15 | 0.279 | |
| Sensitivity 99% | 0.271 | 0.152 | 0.261 | 0.228 | 0.133 | 0.22 | 0.302 | 0.155 | 0.291 | |
| Specificity 90% | 0.902 | 0.903 | 0.903 | 0.906 | 0.906 | 0.907 | 0.679 | 0.896 | 0.902 | |
| Specificity 90% | 0.904 | 0.903 | 0.904 | 0.907 | 0.906 | 0.907 | 0.902 | 0.901 | 0.902 | |
| Specificity 90% | 0.904 | 0.904 | 0.904 | 0.908 | 0.907 | 0.908 | 0.902 | 0.901 | 0.902 | |
| Specificity 90% | 0.905 | 0.904 | 0.904 | 0.908 | 0.907 | 0.908 | 0.902 | 0.901 | 0.902 | |
| Specificity 90% | 0.905 | 0.904 | 0.904 | 0.908 | 0.907 | 0.908 | 0.903 | 0.902 | 0.902 | |
| Specificity 99% | 0.988 | 0.991 | 0.992 | 0.99 | 0.992 | 0.993 | 0.982 | 0.99 | 0.991 | |
| Specificity 99% | 0.991 | 0.991 | 0.992 | 0.992 | 0.992 | 0.993 | 0.99 | 0.99 | 0.991 | |
| Specificity 99% | 0.992 | 0.991 | 0.992 | 0.993 | 0.992 | 0.993 | 0.991 | 0.991 | 0.991 | |
| Specificity 99% | 0.992 | 0.991 | 0.992 | 0.994 | 0.992 | 0.994 | 0.991 | 0.991 | 0.991 | |
| Specificity 99% | 0.992 | 0.991 | 0.992 | 0.994 | 0.992 | 0.994 | 0.991 | 0.991 | 0.991 | |
| PPV 90% | 0.027 | 0.043 | 0.049 | 0.064 | 0.08 | 0.088 | 0.009 | 0.02 | 0.027 | |
| PPV 90% | 0.038 | 0.044 | 0.05 | 0.08 | 0.081 | 0.09 | 0.015 | 0.022 | 0.027 | |
| PPV 90% | 0.053 | 0.046 | 0.052 | 0.096 | 0.084 | 0.093 | 0.028 | 0.023 | 0.028 | |
| PPV 90% | 0.055 | 0.047 | 0.053 | 0.098 | 0.086 | 0.094 | 0.029 | 0.024 | 0.029 | |
| PPV 90% | 0.056 | 0.048 | 0.054 | 0.099 | 0.087 | 0.095 | 0.031 | 0.025 | 0.03 | |
| PPV 99% | 0.081 | 0.118 | 0.204 | 0.192 | 0.21 | 0.327 | 0.028 | 0.054 | 0.114 | |
| PPV 99% | 0.135 | 0.128 | 0.213 | 0.269 | 0.221 | 0.34 | 0.053 | 0.061 | 0.121 | |
| PPV 99% | 0.221 | 0.136 | 0.228 | 0.362 | 0.229 | 0.36 | 0.121 | 0.07 | 0.132 | |
| PPV 99% | 0.235 | 0.138 | 0.232 | 0.375 | 0.233 | 0.367 | 0.134 | 0.071 | 0.136 | |
| PPV 99% | 0.248 | 0.14 | 0.239 | 0.387 | 0.234 | 0.375 | 0.147 | 0.073 | 0.141 | |
| F0.5 90% | 0.033 | 0.052 | 0.06 | 0.076 | 0.096 | 0.105 | 0.011 | 0.024 | 0.033 | |
| F0.5 90% | 0.046 | 0.054 | 0.061 | 0.095 | 0.097 | 0.108 | 0.019 | 0.027 | 0.034 | |
| F0.5 90% | 0.064 | 0.056 | 0.064 | 0.115 | 0.1 | 0.112 | 0.034 | 0.029 | 0.035 | |
| F0.5 90% | 0.067 | 0.057 | 0.065 | 0.117 | 0.103 | 0.113 | 0.036 | 0.029 | 0.036 | |
| F0.5 90% | 0.069 | 0.059 | 0.066 | 0.119 | 0.104 | 0.114 | 0.038 | 0.031 | 0.036 | |
| F0.5 95% | 0.047 | 0.071 | 0.093 | 0.107 | 0.126 | 0.156 | 0.016 | 0.033 | 0.052 | |
| F0.5 95% | 0.069 | 0.073 | 0.096 | 0.139 | 0.129 | 0.161 | 0.028 | 0.037 | 0.054 | |
| F0.5 95% | 0.1 | 0.078 | 0.101 | 0.172 | 0.134 | 0.168 | 0.054 | 0.04 | 0.057 | |
| F0.5 95% | 0.105 | 0.079 | 0.102 | 0.177 | 0.136 | 0.171 | 0.058 | 0.041 | 0.058 | |
| F0.5 95% | 0.108 | 0.081 | 0.104 | 0.18 | 0.137 | 0.173 | 0.062 | 0.043 | 0.059 | |
| F0.5 99% | 0.085 | 0.12 | 0.207 | 0.175 | 0.181 | 0.284 | 0.033 | 0.06 | 0.127 | |
| F0.5 99% | 0.139 | 0.13 | 0.216 | 0.238 | 0.192 | 0.296 | 0.06 | 0.068 | 0.135 | |
| F0.5 99% | 0.225 | 0.138 | 0.232 | 0.318 | 0.199 | 0.315 | 0.135 | 0.078 | 0.147 | |
| F0.5 99% | 0.239 | 0.141 | 0.236 | 0.329 | 0.202 | 0.322 | 0.149 | 0.079 | 0.151 | |
| F0.5 99% | 0.252 | 0.142 | 0.243 | 0.34 | 0.203 | 0.328 | 0.164 | 0.082 | 0.157 | |
| F1 90% | 0.049 | 0.078 | 0.09 | 0.109 | 0.136 | 0.15 | 0.018 | 0.038 | 0.051 | |
| F1 90% | 0.069 | 0.081 | 0.092 | 0.136 | 0.137 | 0.153 | 0.029 | 0.042 | 0.052 | |
| F1 90% | 0.097 | 0.084 | 0.096 | 0.164 | 0.143 | 0.159 | 0.052 | 0.045 | 0.054 | |
| F1 90% | 0.1 | 0.086 | 0.097 | 0.167 | 0.146 | 0.161 | 0.056 | 0.045 | 0.055 | |
| F1 90% | 0.103 | 0.088 | 0.098 | 0.17 | 0.148 | 0.163 | 0.059 | 0.047 | 0.056 | |
| F1 95% | 0.066 | 0.1 | 0.132 | 0.139 | 0.162 | 0.203 | 0.024 | 0.05 | 0.078 | |
| F1 95% | 0.097 | 0.104 | 0.135 | 0.179 | 0.166 | 0.209 | 0.041 | 0.056 | 0.08 | |
| F1 95% | 0.142 | 0.11 | 0.142 | 0.223 | 0.173 | 0.218 | 0.081 | 0.061 | 0.085 | |
| F1 95% | 0.148 | 0.111 | 0.144 | 0.229 | 0.175 | 0.221 | 0.087 | 0.061 | 0.086 | |
| F1 95% | 0.153 | 0.114 | 0.147 | 0.233 | 0.177 | 0.224 | 0.092 | 0.064 | 0.088 | |
| F1 99% | 0.094 | 0.123 | 0.211 | 0.156 | 0.15 | 0.239 | 0.044 | 0.071 | 0.153 | |
| F1 99% | 0.144 | 0.133 | 0.221 | 0.204 | 0.16 | 0.249 | 0.073 | 0.082 | 0.163 | |
| F1 99% | 0.231 | 0.141 | 0.238 | 0.268 | 0.166 | 0.266 | 0.162 | 0.095 | 0.177 | |
| F1 99% | 0.245 | 0.144 | 0.242 | 0.278 | 0.169 | 0.272 | 0.18 | 0.096 | 0.182 | |
| F1 99% | 0.259 | 0.146 | 0.25 | 0.287 | 0.17 | 0.277 | 0.198 | 0.099 | 0.19 | |
| NPV 90% | 0.993 | 0.995 | 0.995 | 0.988 | 0.989 | 0.991 | 0.996 | 0.997 | 0.998 | |
| NPV 90% | 0.994 | 0.995 | 0.995 | 0.99 | 0.989 | 0.991 | 0.996 | 0.997 | 0.998 | |
| NPV 90% | 0.996 | 0.995 | 0.996 | 0.992 | 0.99 | 0.991 | 0.998 | 0.997 | 0.998 | |
| NPV 90% | 0.996 | 0.995 | 0.996 | 0.992 | 0.99 | 0.992 | 0.998 | 0.997 | 0.998 | |
| NPV 90% | 0.996 | 0.995 | 0.996 | 0.992 | 0.99 | 0.992 | 0.998 | 0.998 | 0.998 | |
| NPV 95% | 0.992 | 0.993 | 0.994 | 0.987 | 0.987 | 0.989 | 0.996 | 0.997 | 0.997 | |
| NPV 95% | 0.993 | 0.994 | 0.995 | 0.988 | 0.987 | 0.989 | 0.996 | 0.997 | 0.997 | |
| NPV 95% | 0.995 | 0.994 | 0.995 | 0.99 | 0.988 | 0.99 | 0.997 | 0.997 | 0.997 | |
| NPV 95% | 0.995 | 0.994 | 0.995 | 0.99 | 0.988 | 0.99 | 0.997 | 0.997 | 0.997 | |
| NPV 95% | 0.995 | 0.994 | 0.995 | 0.99 | 0.988 | 0.99 | 0.998 | 0.997 | 0.997 | |
| NPV 99% | 0.992 | 0.992 | 0.993 | 0.985 | 0.984 | 0.986 | 0.996 | 0.996 | 0.996 | |
| NPV 99% | 0.992 | 0.992 | 0.993 | 0.986 | 0.984 | 0.986 | 0.996 | 0.996 | 0.996 | |
| NPV 99% | 0.993 | 0.992 | 0.993 | 0.986 | 0.985 | 0.986 | 0.996 | 0.996 | 0.996 | |
| NPV 99% | 0.993 | 0.992 | 0.993 | 0.987 | 0.985 | 0.986 | 0.996 | 0.996 | 0.996 | |
| NPV 99% | 0.993 | 0.992 | 0.993 | 0.987 | 0.985 | 0.987 | 0.997 | 0.996 | 0.997 | |
| Accuracy 90% | 0.897 | 0.899 | 0.9 | 0.897 | 0.898 | 0.9 | 0.677 | 0.894 | 0.9 | |
| Accuracy 90% | 0.899 | 0.899 | 0.901 | 0.9 | 0.898 | 0.901 | 0.9 | 0.899 | 0.9 | |
| Accuracy 90% | 0.901 | 0.9 | 0.901 | 0.902 | 0.899 | 0.902 | 0.901 | 0.9 | 0.901 | |
| Accuracy 90% | 0.902 | 0.9 | 0.901 | 0.903 | 0.9 | 0.902 | 0.901 | 0.9 | 0.901 | |
| Accuracy 90% | 0.902 | 0.9 | 0.902 | 0.903 | 0.9 | 0.902 | 0.901 | 0.9 | 0.901 | |
| Accuracy 95% | 0.944 | 0.946 | 0.948 | 0.943 | 0.943 | 0.946 | 0.756 | 0.944 | 0.949 | |
| Accuracy 95% | 0.947 | 0.947 | 0.949 | 0.945 | 0.944 | 0.947 | 0.948 | 0.948 | 0.949 | |
| Accuracy 95% | 0.949 | 0.947 | 0.949 | 0.948 | 0.944 | 0.948 | 0.95 | 0.948 | 0.95 | |
| Accuracy 95% | 0.95 | 0.947 | 0.949 | 0.948 | 0.944 | 0.948 | 0.95 | 0.948 | 0.95 | |
| Accuracy 95% | 0.95 | 0.948 | 0.95 | 0.949 | 0.944 | 0.948 | 0.95 | 0.949 | 0.95 | |
| Accuracy 99% | 0.98 | 0.983 | 0.985 | 0.976 | 0.977 | 0.979 | 0.978 | 0.986 | 0.987 | |
| Accuracy 99% | 0.983 | 0.983 | 0.985 | 0.978 | 0.977 | 0.98 | 0.986 | 0.986 | 0.988 | |
| Accuracy 99% | 0.985 | 0.984 | 0.985 | 0.98 | 0.977 | 0.98 | 0.988 | 0.987 | 0.988 | |
| Accuracy 99% | 0.986 | 0.984 | 0.985 | 0.98 | 0.977 | 0.98 | 0.988 | 0.987 | 0.988 | |
| Accuracy 99% | 0.986 | 0.984 | 0.986 | 0.981 | 0.977 | 0.98 | 0.988 | 0.987 | 0.988 | |
| Brier score | 0.013 | 0.009 | 0.009 | 0.019 | 0.017 | 0.015 | 0.015 | 0.009 | 0.005 | |
| Brier score | 0.01 | 0.009 | 0.008 | 0.017 | 0.017 | 0.015 | 0.006 | 0.011 | 0.005 | |
| Brier score | 0.008 | 0.009 | 0.008 | 0.015 | 0.016 | 0.015 | 0.005 | 0.01 | 0.005 | |
| Brier score | 0.008 | 0.009 | 0.008 | 0.015 | 0.016 | 0.015 | 0.005 | 0.01 | 0.005 | |
| Brier score | 0.008 | 0.009 | 0.008 | 0.015 | 0.016 | 0.015 | 0.004 | 0.005 | 0.004 | |
[…]
S1.3 Behavior of AUC
In Figure S1, we show bias and coverage for estimating AUC versus training set size. The rows are the event rates, while the columns are the algorithms. We see that the trends in bias and coverage do not depend on event rate, but instead depend on the number of events.
In Figure S2, we show the empirical mean squared error and confidence interval width for estimating AUC.
S1.4 Behavior of other prediction metrics
We present behavior for sensitivity, specificity, and PPV at the 90th and 99th percentiles of predicted risk; accuracy, NPV, F1, and F0.5 at the 90th, 95th, and 99th percentiles of predicted risk; and the Brier score in Tables S4–S22.
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.045 | 0.065 | 0.947 | 1.131 | ||
| GLM | 46 | -0.02 | 0.007 | 0.940 | 0.307 | ||
| GLM | 230 | -0.008 | 0.001 | 0.944 | 0.136 | ||
| GLM | 460 | -0.004 | 0.001 | 0.948 | 0.090 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.944 | 0.028 | ||
| GLM | 46 | -0.013 | 0.008 | 0.959 | 0.351 | ||
| GLM | 92 | -0.019 | 0.004 | 0.938 | 0.235 | ||
| GLM | 460 | -0.005 | 0.001 | 0.945 | 0.099 | ||
| GLM | 919 | -0.001 | 0.001 | 0.952 | 0.065 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.954 | 0.020 | ||
| GLM | 92 | -0.014 | 0.004 | 0.947 | 0.243 | ||
| GLM | 184 | -0.011 | 0.002 | 0.946 | 0.172 | ||
| GLM | 919 | -0.002 | 0.001 | 0.941 | 0.067 | ||
| GLM | 1837 | -0.001 | 0.001 | 0.952 | 0.047 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.945 | 0.015 | ||
| RF | 23 | -0.006 | 0.014 | 0.951 | 0.487 | ||
| RF | 46 | 0.001 | 0.008 | 0.947 | 0.319 | ||
| RF | 230 | 0.002 | 0.002 | 0.950 | 0.134 | ||
| RF | 460 | 0.002 | 0.002 | 0.948 | 0.097 | ||
| RF | 4591 | 0.001 | 0.001 | 0.946 | 0.028 | ||
| RF | 46 | -0.002 | 0.006 | 0.952 | 0.317 | ||
| RF | 92 | -0.002 | 0.003 | 0.950 | 0.210 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.089 | ||
| RF | 919 | -0.001 | 0.001 | 0.952 | 0.064 | ||
| RF | 9181 | 0.001 | 0.001 | 0.944 | 0.019 | ||
| RF | 92 | -0.002 | 0.003 | 0.953 | 0.216 | ||
| RF | 184 | 0.001 | 0.002 | 0.949 | 0.149 | ||
| RF | 919 | 0.001 | 0.001 | 0.950 | 0.063 | ||
| RF | 1837 | -0.002 | 0.001 | 0.939 | 0.045 | ||
| RF | 18361 | 0.001 | 0.001 | 0.952 | 0.015 | ||
| Ridge | 23 | 0.001 | 0.012 | 0.950 | 0.426 | ||
| Ridge | 46 | -0.002 | 0.005 | 0.946 | 0.287 | ||
| Ridge | 230 | -0.003 | 0.001 | 0.952 | 0.128 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.088 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.952 | 0.028 | ||
| Ridge | 46 | 0.001 | 0.006 | 0.947 | 0.301 | ||
| Ridge | 92 | -0.001 | 0.003 | 0.952 | 0.206 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.950 | 0.091 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.063 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.952 | 0.020 | ||
| Ridge | 92 | -0.003 | 0.003 | 0.947 | 0.214 | ||
| Ridge | 184 | -0.002 | 0.002 | 0.946 | 0.152 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.067 | ||
| Ridge | 1837 | -0.001 | 0.001 | 0.950 | 0.046 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.953 | 0.015 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.009 | 0.005 | 0.958 | 0.301 | ||
| GLM | 46 | -0.008 | 0.003 | 0.952 | 0.193 | ||
| GLM | 230 | -0.006 | 0.001 | 0.946 | 0.116 | ||
| GLM | 460 | -0.003 | 0.001 | 0.949 | 0.080 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.954 | 0.024 | ||
| GLM | 46 | -0.003 | 0.003 | 0.954 | 0.205 | ||
| GLM | 92 | -0.007 | 0.002 | 0.949 | 0.161 | ||
| GLM | 460 | -0.003 | 0.001 | 0.945 | 0.080 | ||
| GLM | 919 | -0.001 | 0.001 | 0.945 | 0.054 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.017 | ||
| GLM | 92 | -0.004 | 0.002 | 0.954 | 0.154 | ||
| GLM | 184 | -0.004 | 0.001 | 0.950 | 0.117 | ||
| GLM | 919 | -0.002 | 0.001 | 0.946 | 0.050 | ||
| GLM | 1837 | -0.001 | 0.001 | 0.946 | 0.034 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.951 | 0.010 | ||
| RF | 23 | -0.001 | 0.005 | 0.958 | 0.317 | ||
| RF | 46 | 0.001 | 0.003 | 0.966 | 0.234 | ||
| RF | 230 | 0.001 | 0.001 | 0.957 | 0.097 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.068 | ||
| RF | 4591 | 0.001 | 0.001 | 0.949 | 0.020 | ||
| RF | 46 | -0.002 | 0.003 | 0.960 | 0.232 | ||
| RF | 92 | -0.002 | 0.002 | 0.948 | 0.160 | ||
| RF | 460 | 0.001 | 0.001 | 0.946 | 0.063 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.044 | ||
| RF | 9181 | 0.001 | 0.001 | 0.945 | 0.014 | ||
| RF | 92 | 0.001 | 0.001 | 0.956 | 0.148 | ||
| RF | 184 | -0.001 | 0.001 | 0.950 | 0.102 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.041 | ||
| RF | 1837 | 0.001 | 0.001 | 0.953 | 0.028 | ||
| RF | 18361 | 0.001 | 0.001 | 0.952 | 0.009 | ||
| Ridge | 23 | -0.004 | 0.008 | 0.962 | 0.369 | ||
| Ridge | 46 | -0.003 | 0.004 | 0.952 | 0.257 | ||
| Ridge | 230 | -0.002 | 0.001 | 0.951 | 0.112 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.948 | 0.080 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.950 | 0.025 | ||
| Ridge | 46 | -0.001 | 0.004 | 0.953 | 0.244 | ||
| Ridge | 92 | -0.001 | 0.002 | 0.952 | 0.173 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.076 | ||
| Ridge | 919 | -0.001 | 0.001 | 0.951 | 0.054 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.952 | 0.016 | ||
| Ridge | 92 | -0.003 | 0.002 | 0.952 | 0.159 | ||
| Ridge | 184 | -0.001 | 0.001 | 0.955 | 0.113 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.953 | 0.049 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.947 | 0.034 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.952 | 0.011 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.048 | 0.075 | 0.952 | 1.215 | ||
| GLM | 46 | 0.001 | 0.001 | 0.998 | 0.034 | ||
| GLM | 230 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| GLM | 460 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.950 | 0.000 | ||
| GLM | 46 | -0.004 | 0.001 | 0.980 | 0.166 | ||
| GLM | 92 | 0.001 | 0.001 | 0.949 | 0.007 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| GLM | 919 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.928 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.947 | 0.011 | ||
| GLM | 184 | 0.001 | 0.001 | 0.949 | 0.007 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.935 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.937 | 0.011 | ||
| RF | 46 | 0.001 | 0.001 | 0.949 | 0.006 | ||
| RF | 230 | 0.001 | 0.001 | 0.953 | 0.002 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 4591 | 0.001 | 0.001 | 0.946 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.946 | 0.008 | ||
| RF | 92 | 0.001 | 0.001 | 0.948 | 0.005 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.955 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.952 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.009 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.006 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.945 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.945 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.946 | 0.006 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.939 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.948 | 0.006 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.944 | 0.005 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.954 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.944 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.952 | 0.007 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.952 | 0.005 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.944 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.927 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.003 | 0.001 | 0.900 | 0.019 | ||
| GLM | 46 | 0.001 | 0.001 | 0.932 | 0.003 | ||
| GLM | 230 | 0.001 | 0.001 | 0.940 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.917 | 0.000 | ||
| GLM | 46 | -0.001 | 0.001 | 0.936 | 0.007 | ||
| GLM | 92 | 0.001 | 0.001 | 0.944 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.938 | 0.006 | ||
| GLM | 184 | 0.001 | 0.001 | 0.943 | 0.003 | ||
| GLM | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.897 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.786 | 0.004 | ||
| RF | 46 | 0.001 | 0.001 | 0.942 | 0.002 | ||
| RF | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.944 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.935 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.950 | 0.004 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.002 | ||
| RF | 460 | 0.001 | 0.001 | 0.954 | 0.001 | ||
| RF | 919 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.925 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.003 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 1837 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| RF | 18361 | 0.001 | 0.001 | 0.961 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.942 | 0.003 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.955 | 0.002 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.861 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.944 | 0.003 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.958 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.941 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.922 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.942 | 0.003 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.941 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.888 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.048 | 0.074 | 0.952 | 1.204 | ||
| GLM | 46 | 0.001 | 0.001 | 0.998 | 0.033 | ||
| GLM | 230 | 0.001 | 0.001 | 0.950 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.943 | 0.000 | ||
| GLM | 46 | -0.004 | 0.001 | 0.981 | 0.163 | ||
| GLM | 92 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| GLM | 460 | 0.001 | 0.001 | 0.946 | 0.003 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.930 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.953 | 0.012 | ||
| GLM | 184 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.003 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.943 | 0.001 | ||
| RF | 23 | 0.001 | 0.001 | 0.937 | 0.011 | ||
| RF | 46 | 0.001 | 0.001 | 0.954 | 0.006 | ||
| RF | 230 | 0.001 | 0.001 | 0.948 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.944 | 0.002 | ||
| RF | 4591 | 0.001 | 0.001 | 0.955 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.951 | 0.009 | ||
| RF | 92 | 0.001 | 0.001 | 0.956 | 0.006 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| RF | 9181 | 0.001 | 0.001 | 0.906 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.952 | 0.010 | ||
| RF | 184 | 0.001 | 0.001 | 0.951 | 0.007 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.003 | ||
| RF | 1837 | 0.001 | 0.001 | 0.941 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.945 | 0.001 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.952 | 0.007 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.948 | 0.005 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.917 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.945 | 0.007 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.947 | 0.005 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.953 | 0.002 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.957 | 0.009 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.950 | 0.006 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.946 | 0.003 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.947 | 0.001 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.039 | 0.064 | 0.917 | 1.188 | ||
| GLM | 46 | 0.001 | 0.001 | 0.999 | 0.028 | ||
| GLM | 230 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| GLM | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.945 | 0.000 | ||
| GLM | 46 | -0.003 | 0.001 | 0.983 | 0.142 | ||
| GLM | 92 | 0.001 | 0.001 | 0.951 | 0.006 | ||
| GLM | 460 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| GLM | 919 | 0.001 | 0.001 | 0.942 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.956 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.941 | 0.009 | ||
| GLM | 184 | 0.001 | 0.001 | 0.944 | 0.006 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.941 | 0.001 | ||
| RF | 23 | 0.001 | 0.001 | 0.923 | 0.008 | ||
| RF | 46 | 0.001 | 0.001 | 0.954 | 0.005 | ||
| RF | 230 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| RF | 460 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.944 | 0.008 | ||
| RF | 92 | 0.001 | 0.001 | 0.951 | 0.005 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.947 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.946 | 0.009 | ||
| RF | 184 | 0.001 | 0.001 | 0.946 | 0.006 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.003 | ||
| RF | 1837 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.955 | 0.005 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.946 | 0.004 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.947 | 0.006 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.948 | 0.005 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.943 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.945 | 0.006 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.945 | 0.002 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.936 | 0.001 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.003 | 0.001 | 0.890 | 0.017 | ||
| GLM | 46 | 0.001 | 0.001 | 0.930 | 0.004 | ||
| GLM | 230 | 0.001 | 0.001 | 0.938 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.954 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.941 | 0.000 | ||
| GLM | 46 | -0.001 | 0.001 | 0.921 | 0.006 | ||
| GLM | 92 | 0.001 | 0.001 | 0.942 | 0.004 | ||
| GLM | 460 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| GLM | 919 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.938 | 0.008 | ||
| GLM | 184 | 0.001 | 0.001 | 0.952 | 0.005 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.925 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.849 | 0.005 | ||
| RF | 46 | 0.001 | 0.001 | 0.961 | 0.004 | ||
| RF | 230 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.942 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.887 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.953 | 0.006 | ||
| RF | 92 | 0.001 | 0.001 | 0.954 | 0.004 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.943 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.949 | 0.008 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.005 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.949 | 0.004 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.958 | 0.003 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.922 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.956 | 0.005 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.947 | 0.004 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.945 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.933 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.956 | 0.007 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.946 | 0.005 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.956 | 0.002 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.930 | 0.001 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.950 | 0.021 | ||
| GLM | 46 | 0.001 | 0.001 | 0.950 | 0.020 | ||
| GLM | 230 | 0.001 | 0.001 | 0.951 | 0.010 | ||
| GLM | 460 | 0.001 | 0.001 | 0.951 | 0.007 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| GLM | 46 | -0.001 | 0.001 | 0.952 | 0.037 | ||
| GLM | 92 | -0.001 | 0.001 | 0.945 | 0.029 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.014 | ||
| GLM | 919 | 0.001 | 0.001 | 0.952 | 0.009 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.952 | 0.003 | ||
| GLM | 92 | -0.002 | 0.001 | 0.948 | 0.055 | ||
| GLM | 184 | -0.001 | 0.001 | 0.946 | 0.041 | ||
| GLM | 919 | 0.001 | 0.001 | 0.948 | 0.017 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.012 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.953 | 0.004 | ||
| RF | 23 | 0.001 | 0.001 | 0.954 | 0.028 | ||
| RF | 46 | 0.001 | 0.001 | 0.942 | 0.020 | ||
| RF | 230 | 0.001 | 0.001 | 0.952 | 0.009 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.006 | ||
| RF | 4591 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 46 | 0.001 | 0.001 | 0.955 | 0.039 | ||
| RF | 92 | 0.001 | 0.001 | 0.951 | 0.027 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.012 | ||
| RF | 919 | 0.001 | 0.001 | 0.942 | 0.008 | ||
| RF | 9181 | 0.001 | 0.001 | 0.943 | 0.003 | ||
| RF | 92 | 0.001 | 0.001 | 0.954 | 0.051 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.035 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.015 | ||
| RF | 1837 | 0.001 | 0.001 | 0.951 | 0.011 | ||
| RF | 18361 | 0.001 | 0.001 | 0.951 | 0.004 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.949 | 0.029 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.952 | 0.021 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.944 | 0.010 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.007 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.955 | 0.002 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.955 | 0.040 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.955 | 0.029 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.954 | 0.013 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.951 | 0.009 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.948 | 0.003 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.948 | 0.052 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.948 | 0.037 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.017 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.947 | 0.011 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.952 | 0.004 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.001 | 0.001 | 0.960 | 0.079 | ||
| GLM | 46 | -0.004 | 0.001 | 0.959 | 0.102 | ||
| GLM | 230 | -0.002 | 0.001 | 0.952 | 0.066 | ||
| GLM | 460 | 0.001 | 0.001 | 0.952 | 0.046 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.948 | 0.014 | ||
| GLM | 46 | -0.005 | 0.002 | 0.964 | 0.167 | ||
| GLM | 92 | -0.006 | 0.001 | 0.951 | 0.164 | ||
| GLM | 460 | -0.002 | 0.001 | 0.944 | 0.085 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.057 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.018 | ||
| GLM | 92 | -0.008 | 0.004 | 0.963 | 0.263 | ||
| GLM | 184 | -0.006 | 0.003 | 0.948 | 0.216 | ||
| GLM | 919 | -0.003 | 0.001 | 0.950 | 0.096 | ||
| GLM | 1837 | -0.001 | 0.001 | 0.952 | 0.066 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.951 | 0.020 | ||
| RF | 23 | 0.002 | 0.002 | 0.954 | 0.187 | ||
| RF | 46 | 0.001 | 0.001 | 0.965 | 0.123 | ||
| RF | 230 | 0.001 | 0.001 | 0.958 | 0.050 | ||
| RF | 460 | 0.001 | 0.001 | 0.947 | 0.035 | ||
| RF | 4591 | 0.001 | 0.001 | 0.940 | 0.010 | ||
| RF | 46 | 0.002 | 0.003 | 0.958 | 0.252 | ||
| RF | 92 | 0.001 | 0.001 | 0.950 | 0.165 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.063 | ||
| RF | 919 | 0.001 | 0.001 | 0.951 | 0.044 | ||
| RF | 9181 | 0.001 | 0.001 | 0.948 | 0.014 | ||
| RF | 92 | 0.002 | 0.005 | 0.956 | 0.302 | ||
| RF | 184 | 0.001 | 0.002 | 0.947 | 0.199 | ||
| RF | 919 | 0.001 | 0.001 | 0.954 | 0.076 | ||
| RF | 1837 | 0.001 | 0.001 | 0.956 | 0.053 | ||
| RF | 18361 | 0.001 | 0.001 | 0.958 | 0.017 | ||
| Ridge | 23 | 0.002 | 0.003 | 0.948 | 0.224 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.948 | 0.152 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.950 | 0.065 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.946 | 0.046 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.944 | 0.014 | ||
| Ridge | 46 | 0.005 | 0.005 | 0.948 | 0.285 | ||
| Ridge | 92 | 0.004 | 0.002 | 0.946 | 0.193 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.950 | 0.081 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.951 | 0.058 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.948 | 0.017 | ||
| Ridge | 92 | 0.002 | 0.006 | 0.954 | 0.332 | ||
| Ridge | 184 | 0.004 | 0.003 | 0.957 | 0.219 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.951 | 0.092 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.946 | 0.064 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.952 | 0.020 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.958 | 0.004 | ||
| GLM | 46 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| GLM | 230 | 0.001 | 0.001 | 0.928 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.932 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.892 | 0.000 | ||
| GLM | 46 | 0.001 | 0.001 | 0.939 | 0.005 | ||
| GLM | 92 | 0.001 | 0.001 | 0.942 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.944 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.911 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.949 | 0.006 | ||
| GLM | 184 | 0.001 | 0.001 | 0.945 | 0.004 | ||
| GLM | 919 | 0.001 | 0.001 | 0.942 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.917 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.939 | 0.003 | ||
| RF | 46 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| RF | 230 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.937 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.952 | 0.004 | ||
| RF | 92 | 0.001 | 0.001 | 0.947 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| RF | 919 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.885 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.006 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.004 | ||
| RF | 919 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.955 | 0.001 | ||
| RF | 18361 | 0.001 | 0.001 | 0.936 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.956 | 0.003 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.943 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.943 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.956 | 0.004 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.947 | 0.003 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.947 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.956 | 0.006 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.949 | 0.004 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.955 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.939 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.937 | 0.004 | ||
| GLM | 46 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| GLM | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.943 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.858 | 0.000 | ||
| GLM | 46 | 0.001 | 0.001 | 0.947 | 0.005 | ||
| GLM | 92 | 0.001 | 0.001 | 0.944 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.942 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.920 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.946 | 0.006 | ||
| GLM | 184 | 0.001 | 0.001 | 0.945 | 0.004 | ||
| GLM | 919 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.922 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.947 | 0.003 | ||
| RF | 46 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| RF | 230 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.939 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.944 | 0.005 | ||
| RF | 92 | 0.001 | 0.001 | 0.949 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.944 | 0.001 | ||
| RF | 919 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.937 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.957 | 0.007 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.005 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| RF | 18361 | 0.001 | 0.001 | 0.956 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.945 | 0.003 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.941 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.856 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.949 | 0.004 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.947 | 0.003 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.942 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.926 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.950 | 0.006 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.946 | 0.004 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.917 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.950 | 0.004 | ||
| GLM | 46 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| GLM | 230 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.928 | 0.000 | ||
| GLM | 46 | 0.001 | 0.001 | 0.954 | 0.005 | ||
| GLM | 92 | 0.001 | 0.001 | 0.956 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.954 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.926 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.946 | 0.007 | ||
| GLM | 184 | 0.001 | 0.001 | 0.941 | 0.004 | ||
| GLM | 919 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.930 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| RF | 46 | 0.001 | 0.001 | 0.955 | 0.003 | ||
| RF | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.938 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.950 | 0.005 | ||
| RF | 92 | 0.001 | 0.001 | 0.954 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.924 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.951 | 0.007 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.005 | ||
| RF | 919 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.958 | 0.003 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.937 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.960 | 0.005 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.953 | 0.003 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.937 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.954 | 0.007 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.946 | 0.005 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.953 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.943 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.950 | 0.039 | ||
| GLM | 46 | -0.002 | 0.001 | 0.944 | 0.037 | ||
| GLM | 230 | 0.001 | 0.001 | 0.951 | 0.019 | ||
| GLM | 460 | 0.001 | 0.001 | 0.951 | 0.013 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| GLM | 46 | -0.002 | 0.001 | 0.951 | 0.065 | ||
| GLM | 92 | -0.003 | 0.001 | 0.946 | 0.052 | ||
| GLM | 460 | 0.001 | 0.001 | 0.949 | 0.025 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.016 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.955 | 0.005 | ||
| GLM | 92 | -0.004 | 0.001 | 0.948 | 0.089 | ||
| GLM | 184 | -0.003 | 0.001 | 0.944 | 0.065 | ||
| GLM | 919 | 0.001 | 0.001 | 0.946 | 0.027 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.019 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.957 | 0.006 | ||
| RF | 23 | 0.001 | 0.001 | 0.952 | 0.052 | ||
| RF | 46 | 0.001 | 0.001 | 0.942 | 0.037 | ||
| RF | 230 | 0.001 | 0.001 | 0.953 | 0.016 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.012 | ||
| RF | 4591 | 0.001 | 0.001 | 0.950 | 0.003 | ||
| RF | 46 | 0.001 | 0.001 | 0.955 | 0.068 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.047 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.020 | ||
| RF | 919 | 0.001 | 0.001 | 0.941 | 0.015 | ||
| RF | 9181 | 0.001 | 0.001 | 0.945 | 0.005 | ||
| RF | 92 | 0.001 | 0.001 | 0.955 | 0.080 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.055 | ||
| RF | 919 | 0.001 | 0.001 | 0.947 | 0.024 | ||
| RF | 1837 | 0.001 | 0.001 | 0.946 | 0.017 | ||
| RF | 18361 | 0.001 | 0.001 | 0.949 | 0.006 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.948 | 0.054 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.952 | 0.039 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.945 | 0.018 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.948 | 0.012 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.954 | 0.004 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.956 | 0.069 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.954 | 0.050 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.953 | 0.023 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.016 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.952 | 0.005 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.946 | 0.082 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.950 | 0.059 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.026 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.948 | 0.018 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.951 | 0.006 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.953 | 0.062 | ||
| GLM | 46 | -0.002 | 0.001 | 0.952 | 0.059 | ||
| GLM | 230 | -0.001 | 0.001 | 0.949 | 0.032 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.022 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.943 | 0.007 | ||
| GLM | 46 | -0.004 | 0.001 | 0.954 | 0.097 | ||
| GLM | 92 | -0.005 | 0.001 | 0.945 | 0.082 | ||
| GLM | 460 | -0.001 | 0.001 | 0.944 | 0.039 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.026 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| GLM | 92 | -0.006 | 0.001 | 0.948 | 0.124 | ||
| GLM | 184 | -0.005 | 0.001 | 0.939 | 0.095 | ||
| GLM | 919 | -0.001 | 0.001 | 0.947 | 0.040 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.951 | 0.028 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.946 | 0.009 | ||
| RF | 23 | 0.001 | 0.001 | 0.956 | 0.084 | ||
| RF | 46 | 0.001 | 0.001 | 0.953 | 0.060 | ||
| RF | 230 | 0.001 | 0.001 | 0.945 | 0.026 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.019 | ||
| RF | 4591 | 0.001 | 0.001 | 0.942 | 0.006 | ||
| RF | 46 | 0.001 | 0.001 | 0.952 | 0.107 | ||
| RF | 92 | 0.001 | 0.001 | 0.948 | 0.074 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.031 | ||
| RF | 919 | 0.001 | 0.001 | 0.951 | 0.022 | ||
| RF | 9181 | 0.001 | 0.001 | 0.948 | 0.007 | ||
| RF | 92 | -0.001 | 0.001 | 0.954 | 0.115 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.078 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.034 | ||
| RF | 1837 | 0.001 | 0.001 | 0.945 | 0.024 | ||
| RF | 18361 | 0.001 | 0.001 | 0.941 | 0.008 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.951 | 0.093 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.951 | 0.067 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.952 | 0.030 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.948 | 0.021 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.953 | 0.007 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.954 | 0.112 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.951 | 0.080 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.948 | 0.036 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.947 | 0.025 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.946 | 0.008 | ||
| Ridge | 92 | -0.002 | 0.001 | 0.950 | 0.121 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.955 | 0.088 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.951 | 0.039 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.948 | 0.026 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.950 | 0.009 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.002 | 0.001 | 0.965 | 0.111 | ||
| GLM | 46 | -0.007 | 0.001 | 0.963 | 0.126 | ||
| GLM | 230 | -0.004 | 0.001 | 0.947 | 0.082 | ||
| GLM | 460 | -0.002 | 0.001 | 0.951 | 0.057 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.952 | 0.017 | ||
| GLM | 46 | -0.008 | 0.002 | 0.964 | 0.167 | ||
| GLM | 92 | -0.01 | 0.001 | 0.950 | 0.153 | ||
| GLM | 460 | -0.004 | 0.001 | 0.937 | 0.079 | ||
| GLM | 919 | -0.002 | 0.001 | 0.944 | 0.053 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.017 | ||
| GLM | 92 | -0.012 | 0.002 | 0.950 | 0.177 | ||
| GLM | 184 | -0.009 | 0.001 | 0.940 | 0.143 | ||
| GLM | 919 | -0.004 | 0.001 | 0.943 | 0.063 | ||
| GLM | 1837 | -0.002 | 0.001 | 0.950 | 0.043 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.956 | 0.013 | ||
| RF | 23 | -0.003 | 0.002 | 0.965 | 0.209 | ||
| RF | 46 | -0.002 | 0.001 | 0.965 | 0.152 | ||
| RF | 230 | 0.001 | 0.001 | 0.959 | 0.064 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.045 | ||
| RF | 4591 | 0.001 | 0.001 | 0.943 | 0.013 | ||
| RF | 46 | -0.006 | 0.003 | 0.962 | 0.219 | ||
| RF | 92 | -0.005 | 0.001 | 0.947 | 0.153 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.061 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.043 | ||
| RF | 9181 | 0.001 | 0.001 | 0.946 | 0.013 | ||
| RF | 92 | -0.006 | 0.002 | 0.953 | 0.182 | ||
| RF | 184 | -0.005 | 0.001 | 0.950 | 0.127 | ||
| RF | 919 | 0.001 | 0.001 | 0.952 | 0.052 | ||
| RF | 1837 | -0.001 | 0.001 | 0.950 | 0.036 | ||
| RF | 18361 | 0.001 | 0.001 | 0.949 | 0.011 | ||
| Ridge | 23 | -0.008 | 0.004 | 0.947 | 0.247 | ||
| Ridge | 46 | -0.005 | 0.002 | 0.947 | 0.177 | ||
| Ridge | 230 | -0.002 | 0.001 | 0.952 | 0.079 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.943 | 0.056 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.942 | 0.018 | ||
| Ridge | 46 | -0.009 | 0.003 | 0.950 | 0.235 | ||
| Ridge | 92 | -0.005 | 0.002 | 0.944 | 0.169 | ||
| Ridge | 460 | -0.002 | 0.001 | 0.950 | 0.075 | ||
| Ridge | 919 | -0.002 | 0.001 | 0.952 | 0.054 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.955 | 0.016 | ||
| Ridge | 92 | -0.011 | 0.003 | 0.946 | 0.196 | ||
| Ridge | 184 | -0.006 | 0.001 | 0.953 | 0.140 | ||
| Ridge | 919 | -0.002 | 0.001 | 0.954 | 0.061 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.945 | 0.043 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.948 | 0.014 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.009 | 0.001 | 0.860 | 0.041 | ||
| GLM | 46 | 0.003 | 0.001 | 0.894 | 0.020 | ||
| GLM | 230 | 0.001 | 0.001 | 0.951 | 0.013 | ||
| GLM | 460 | 0.001 | 0.001 | 0.952 | 0.009 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.950 | 0.003 | ||
| GLM | 46 | 0.008 | 0.001 | 0.863 | 0.037 | ||
| GLM | 92 | -0.001 | 0.001 | 0.950 | 0.034 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.017 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.011 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.954 | 0.004 | ||
| GLM | 92 | 0.001 | 0.001 | 0.954 | 0.060 | ||
| GLM | 184 | -0.002 | 0.001 | 0.946 | 0.048 | ||
| GLM | 919 | 0.001 | 0.001 | 0.947 | 0.020 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.014 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.955 | 0.004 | ||
| RF | 23 | 0.009 | 0.001 | 0.732 | 0.025 | ||
| RF | 46 | 0.002 | 0.001 | 0.942 | 0.022 | ||
| RF | 230 | 0.001 | 0.001 | 0.953 | 0.011 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| RF | 4591 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| RF | 46 | 0.003 | 0.001 | 0.941 | 0.041 | ||
| RF | 92 | 0.001 | 0.001 | 0.952 | 0.032 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.014 | ||
| RF | 919 | 0.001 | 0.001 | 0.942 | 0.010 | ||
| RF | 9181 | 0.001 | 0.001 | 0.943 | 0.003 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.058 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.041 | ||
| RF | 919 | 0.001 | 0.001 | 0.946 | 0.018 | ||
| RF | 1837 | 0.001 | 0.001 | 0.950 | 0.013 | ||
| RF | 18361 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| Ridge | 23 | 0.006 | 0.001 | 0.863 | 0.029 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.946 | 0.024 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.944 | 0.012 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.008 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.957 | 0.003 | ||
| Ridge | 46 | 0.002 | 0.001 | 0.942 | 0.045 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.954 | 0.034 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.954 | 0.016 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.011 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.948 | 0.003 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.949 | 0.060 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.948 | 0.044 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.019 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.947 | 0.013 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.951 | 0.004 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.022 | 0.001 | 0.843 | 0.084 | ||
| GLM | 46 | 0.01 | 0.001 | 0.786 | 0.033 | ||
| GLM | 230 | 0.001 | 0.001 | 0.950 | 0.022 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.015 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.945 | 0.005 | ||
| GLM | 46 | 0.022 | 0.001 | 0.730 | 0.061 | ||
| GLM | 92 | 0.001 | 0.001 | 0.951 | 0.055 | ||
| GLM | 460 | 0.001 | 0.001 | 0.944 | 0.029 | ||
| GLM | 919 | 0.001 | 0.001 | 0.952 | 0.020 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.952 | 0.006 | ||
| GLM | 92 | 0.005 | 0.001 | 0.948 | 0.088 | ||
| GLM | 184 | -0.003 | 0.001 | 0.942 | 0.078 | ||
| GLM | 919 | 0.001 | 0.001 | 0.947 | 0.033 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.024 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.947 | 0.007 | ||
| RF | 23 | 0.025 | 0.001 | 0.420 | 0.043 | ||
| RF | 46 | 0.007 | 0.001 | 0.878 | 0.035 | ||
| RF | 230 | 0.001 | 0.001 | 0.947 | 0.018 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.013 | ||
| RF | 4591 | 0.001 | 0.001 | 0.946 | 0.004 | ||
| RF | 46 | 0.015 | 0.001 | 0.863 | 0.065 | ||
| RF | 92 | 0.002 | 0.001 | 0.948 | 0.051 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.023 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.016 | ||
| RF | 9181 | 0.001 | 0.001 | 0.946 | 0.005 | ||
| RF | 92 | 0.005 | 0.001 | 0.946 | 0.089 | ||
| RF | 184 | 0.001 | 0.001 | 0.952 | 0.064 | ||
| RF | 919 | 0.001 | 0.001 | 0.950 | 0.028 | ||
| RF | 1837 | 0.001 | 0.001 | 0.946 | 0.020 | ||
| RF | 18361 | 0.001 | 0.001 | 0.945 | 0.006 | ||
| Ridge | 23 | 0.018 | 0.001 | 0.747 | 0.051 | ||
| Ridge | 46 | 0.003 | 0.001 | 0.940 | 0.041 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.953 | 0.021 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.015 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.951 | 0.005 | ||
| Ridge | 46 | 0.009 | 0.001 | 0.921 | 0.074 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.950 | 0.058 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.027 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.019 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.945 | 0.006 | ||
| Ridge | 92 | 0.002 | 0.001 | 0.950 | 0.097 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.956 | 0.073 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.954 | 0.032 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.948 | 0.022 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.951 | 0.007 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.096 | 0.012 | 0.606 | 0.195 | ||
| GLM | 46 | 0.067 | 0.005 | 0.431 | 0.116 | ||
| GLM | 230 | -0.003 | 0.001 | 0.947 | 0.071 | ||
| GLM | 460 | -0.001 | 0.001 | 0.953 | 0.050 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.947 | 0.015 | ||
| GLM | 46 | 0.102 | 0.013 | 0.494 | 0.193 | ||
| GLM | 92 | 0.03 | 0.002 | 0.854 | 0.133 | ||
| GLM | 460 | -0.003 | 0.001 | 0.945 | 0.082 | ||
| GLM | 919 | -0.001 | 0.001 | 0.948 | 0.055 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.952 | 0.018 | ||
| GLM | 92 | 0.06 | 0.006 | 0.758 | 0.178 | ||
| GLM | 184 | 0.002 | 0.002 | 0.954 | 0.161 | ||
| GLM | 919 | -0.004 | 0.001 | 0.943 | 0.078 | ||
| GLM | 1837 | -0.002 | 0.001 | 0.950 | 0.053 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.953 | 0.016 | ||
| RF | 23 | 0.186 | 0.041 | 0.364 | 0.309 | ||
| RF | 46 | 0.072 | 0.006 | 0.416 | 0.126 | ||
| RF | 230 | 0.002 | 0.001 | 0.954 | 0.053 | ||
| RF | 460 | 0.001 | 0.001 | 0.946 | 0.038 | ||
| RF | 4591 | 0.001 | 0.001 | 0.942 | 0.011 | ||
| RF | 46 | 0.126 | 0.019 | 0.430 | 0.229 | ||
| RF | 92 | 0.041 | 0.003 | 0.777 | 0.132 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.062 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.043 | ||
| RF | 9181 | 0.001 | 0.001 | 0.944 | 0.013 | ||
| RF | 92 | 0.077 | 0.008 | 0.661 | 0.192 | ||
| RF | 184 | 0.015 | 0.001 | 0.927 | 0.142 | ||
| RF | 919 | 0.001 | 0.001 | 0.953 | 0.063 | ||
| RF | 1837 | -0.001 | 0.001 | 0.951 | 0.044 | ||
| RF | 18361 | 0.001 | 0.001 | 0.953 | 0.014 | ||
| Ridge | 23 | 0.137 | 0.023 | 0.446 | 0.247 | ||
| Ridge | 46 | 0.044 | 0.003 | 0.754 | 0.134 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.951 | 0.069 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.945 | 0.049 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.945 | 0.016 | ||
| Ridge | 46 | 0.09 | 0.011 | 0.618 | 0.210 | ||
| Ridge | 92 | 0.021 | 0.002 | 0.916 | 0.159 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.951 | 0.078 | ||
| Ridge | 919 | -0.001 | 0.001 | 0.952 | 0.056 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.953 | 0.017 | ||
| Ridge | 92 | 0.042 | 0.004 | 0.877 | 0.205 | ||
| Ridge | 184 | 0.002 | 0.002 | 0.957 | 0.164 | ||
| Ridge | 919 | -0.002 | 0.001 | 0.950 | 0.075 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.946 | 0.052 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.949 | 0.017 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.002 | 0.001 | 0.918 | 0.021 | ||
| GLM | 46 | 0.001 | 0.001 | 0.925 | 0.003 | ||
| GLM | 230 | 0.001 | 0.001 | 0.955 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.896 | 0.000 | ||
| GLM | 46 | 0.001 | 0.001 | 0.926 | 0.006 | ||
| GLM | 92 | 0.001 | 0.001 | 0.939 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.947 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.941 | 0.007 | ||
| GLM | 184 | 0.001 | 0.001 | 0.938 | 0.004 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.929 | 0.000 | ||
| RF | 23 | -0.004 | 0.004 | 0.021 | 0.004 | ||
| RF | 46 | -0.006 | 0.006 | 0.000 | 0.003 | ||
| RF | 230 | -0.005 | 0.005 | 0.000 | 0.001 | ||
| RF | 460 | -0.006 | 0.006 | 0.000 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.951 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.949 | 0.005 | ||
| RF | 92 | 0.001 | 0.001 | 0.952 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.954 | 0.007 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.005 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| RF | 18361 | 0.001 | 0.001 | 0.943 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.952 | 0.004 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.951 | 0.003 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.944 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.930 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.955 | 0.005 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.951 | 0.003 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.947 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.955 | 0.006 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.946 | 0.004 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.931 | 0.000 |
S1.5 Determining fixed tuning parameters
To determine the values of fixed tuning parameters, we ran a small pilot simulation study. Here, we again varied and considered the three event rates defined in the main manuscript. In this pilot simulation, we bootstrap sampled outcomes along with covariates; we do not expect that this made a meaningful difference in the behavior of AUC. This investigation consisted of two parts: 1) nested cross-validation for tuning parameter selection (inner folds) and estimating AUC (outer folds); and 2) a small number of simulations, with relatively few replications, to assess if there was a large difference between the cross-validated AUC when the selected fixed tuning parameters were used and the cross-validated AUC obtained when we used 20-fold cross-validation to select the tuning parameter in the sampled training set. We referred to the second part of this fixed tuning parameter investigations as pilot simulations. For ridge regression, we used internal tuning from the R function cv.glmnet. For random forests, we set the number of trees to be 250 for , 100 for , and 50 for and tuned over grids of minimum node size given in Table S23. We assessed how variable the selected tuning parameters were across simulation replicates, and determined that the variability was small enough to justify fixing the tuning parameters as described above.
| Training set size | Grid of minimum node sizes |