### v2

# Behavior of prediction performance metrics with rare events

Abstract
Objective Area under the receiving operator characteristic curve (AUC) is commonly reported alongside prediction models for binary outcomes. Recent articles have raised concerns that AUC might be a misleading measure of prediction performance in the rare event setting. This setting is common since many events of clinical importance are rare. We aimed to determine whether the bias and variance of AUC are driven by the number of events or the event rate. We also investigated the behavior of other commonly used measures of prediction performance, including positive predictive value, accuracy, sensitivity, and specificity.
Study Design and Setting We conducted a simulation study to determine when or whether AUC is unstable in the rare event setting by varying the size of datasets used to train and evaluate prediction models. This plasmode simulation study was based on data from the Mental Health Research Network; the data contained 149 predictors and the outcome of interest, suicide attempt, which had event rate 0.92% in the original dataset.
Results Our results indicate that poor AUC behavior—as measured by empirical bias, variability of cross-validated AUC estimates, and empirical coverage of confidence intervals—is driven by the number of events in a rare-event setting, not event rate. Performance of sensitivity is driven by the number of events, while that of specificity is driven by the number of non-events. Other measures, including positive predictive value and accuracy, depend on the event rate even in large samples.
Conclusion AUC is reliable in the rare event setting provided that the total number of events is moderately large; in our simulations, we observed near zero bias with 1000 events.
Plain Language Summary Predicting self-harm or suicidal behavior is medically important for guiding clinicians in providing care to patients. Several research teams have developed and evaluated suicide risk prediction models based on health records data. Part of evaluating these models is calculating area under the receiver operating characteristic curve (AUC) and other prediction performance metrics. Self-harm and suicide are rare events. Recent research has raised concerns with using AUC in rare-event settings. We aimed to determine whether having a sufficiently large dataset could remove these concerns. In our experiments, we found that AUC can be used without concern in settings with 1000 events or more. Thus, AUC is a valid measure of suicide risk prediction model performance in many large healthcare databases.
Keywords: Prediction, classification, machine learning, area under the receiver operating characteristic curve, model evaluation, rare outcome.
What is new
Key findings
AUC and other binary-outcome prediction performance metrics can be used reliably in rare-outcome settings so long as the minimum class size is large enough.
Multiple prediction performance metrics should be reported in binary outcome settings, along with estimates of variability.
What this adds to what is known?
We provide evidence that the bias and variance of AUC, sensitivity, and specificity are driven by effective sample size (number of events for sensitivity; number of non-events for specificity; and the minimum of these two for AUC). In our simulations, we found that 1000 events was sufficient to have near zero bias for AUC and near nominal confidence interval coverage.
We provide evidence that the bias and variance of other binary outcome prediction performance metrics also depends on the event rate.
What is the implication and what should change now?
Our results imply that in settings with rare events but large sample sizes—for example, predicting suicide attempt using millions of health care visits—AUC and other prediction performance metrics should reliably capture prediction performance. These measures should be used when scientifically relevant.
1 Introduction
Clinical decision-making can be guided by prediction models that identify people at risk for negative health outcomes [1, 2, 3, 4]. Assessing model performance across many metrics is important. Selecting any one particular metric depends on how the model will be used in care [5, 6, 7]. Sensitivity and specificity [8], Brier score [9], and the area under the receiver operating characteristic curve [10, AUC;] are all useful performance metrics when considering binary outcomes. Estimating and reporting prediction model performance in the rare event setting presents several challenges. First, even with a large number of observations, the absolute number of events may be small. In training predictive models, the number of events, not just the total sample size, is of importance [11]. Second, a small event rate impacts the behavior of some performance measures [12, 13].
AUC is often reported in the scientific literature to describe the performance of risk prediction models, as we and others have done when estimating models to predict suicide attempt [14, 15, 4, 16, 17, 18, 19, 20, 21, 22, 23]. In our prior work, the per-visit 90-day suicide attempt rate was 0.62% for visits to mental health specialty clinics and 0.26% for general medical visits with mental health diagnoses; suicide deaths were even more rare, approximately 0.02% and 0.01%, respectively [4]. Recently, several studies have argued that AUC is a misleading and inappropriate measure of prediction model performance in the rare event setting [13, 12, 24, 25, 26]. In practice, when using large sample sizes, even with very rare events, AUC estimates of suicide risk prediction model performance have aligned with other performance metrics; in other words, multiple metrics yield similar conclusions about model adequacy and strength of risk discrimination [14, 18, 23]. This is contrary to concerns raised in recent literature [12, 25]. It is essential for researchers and decision makers to have a clear understanding of the statistical properties of performance measures in rare-outcome settings, including bias (including the potential for optimism) and stability (i.e., estimator variability), as many prediction models are developed to predict rare outcomes. In this paper we examine bias and variance of AUC and other performance metrics in a variety of settings, varying both the sample size and the event rate.
2 Materials and Methods
2.1 Prediction performance metrics: notation and definitions
We assume independent and identically distributed observations drawn from a distribution , with data unit , covariate vector and binary outcome . We further assume a prediction model, , trained on a sample from , such that returns a predicted probability of given . The AUC of is
where and are independent draws from . In other words, AUC is the probability that a randomly sampled non-event has a lower predicted probability than a randomly sampled event. We define other common performance metrics in Table 1.
| Metric | Definition |
| Sensitivity | |
| Specificity | |
| Positive predictive value | |
| Negative predictive value | |
| F0.5 score | |
| F1 score | |
| Classification accuracy | |
| Brier score |
2.2 Simulation methods
We examined the behavior of AUC and the performance metrics in Table 1 in a variety of rare-event settings, with both large and small sample size.
2.2.1 Data
To maintain both the complex outcome-predictor relationships and the complex correlation structure among predictors in real-world data [4, 20], we used a plasmode simulation approach [27]. We obtained a dataset with independent observations by randomly sampling one mental health visit per patient from across seven Mental Health Research Network sites (https://mhresearchnetwork.org). These visits were made by patients aged 13 years or older and occurred between January 1, 2009 and September 30, 2017 [20]. For patients with a suicide attempt (fatal or non-fatal) during this time, we sampled a mental health visit within 90 days; for those without a suicide attempt, a randomly sampled mental health visit was selected. The event rate in , , was 0.92%. We considered 149 predictors, including demographic characteristics, mental health and substance abuse diagnoses (including past self-inflicted injuries or poisonings), other medical diagnoses, past inpatient or emergency mental health care, dispensed mental health medications, the patient health questionnaire (PHQ) 8-item score (PHQ-8) representing depressive symptoms, and the PHQ 9th item response indicating thoughts of death or self-harm [4]. All predictors except patient age in years and PHQ-8 score (0–24) were coded as binary indicators. The PHQ item 9, which has 4 possible responses (0,1,2,3), was coded using 3 binary variables indicating each possible non-zero response.
2.2.2 Simulation scenarios
We used three types of outcome-generating functions for our simulations: logistic regression, ridge regression [28], and random forest models [29]. Outcome-generating models for event rates of 0.92%, 0.46%, and 1.84% were created by sampling 1 million visits from with and without events to match the desired event rate. The tuning parameters used for each outcome-generating model are provided in Table S1; 10-fold cross-validated estimates of performance of these models is presented in Table 2.
| Algorithm | Event rate | AUC | Brier | Percentile | Sens. | Spec. | Acc. | PPV | NPV | F1 | F0.5 |
| 90 | 0.613 | 0.876 | 0.875 | 0.022 | 0.998 | 0.043 | 0.028 | ||||
| 95 | 0.503 | 0.93 | 0.928 | 0.032 | 0.998 | 0.06 | 0.039 | ||||
| 0.005 | 0.806 | 0.004 | 99 | 0.263 | 0.989 | 0.985 | 0.097 | 0.997 | 0.141 | 0.111 | |
| 90 | 0.65 | 0.872 | 0.87 | 0.045 | 0.996 | 0.084 | 0.055 | ||||
| 95 | 0.534 | 0.936 | 0.932 | 0.072 | 0.995 | 0.127 | 0.087 | ||||
| 0.009 | 0.832 | 0.008 | 99 | 0.194 | 0.996 | 0.988 | 0.296 | 0.993 | 0.234 | 0.268 | |
| 90 | 0.671 | 0.881 | 0.877 | 0.095 | 0.993 | 0.167 | 0.115 | ||||
| 95 | 0.536 | 0.945 | 0.937 | 0.154 | 0.991 | 0.239 | 0.18 | ||||
| RF | 0.018 | 0.854 | 0.015 | 99 | 0.144 | 0.998 | 0.982 | 0.568 | 0.984 | 0.229 | 0.357 |
| 90 | 0.642 | 0.902 | 0.901 | 0.029 | 0.998 | 0.056 | 0.036 | ||||
| 95 | 0.517 | 0.952 | 0.95 | 0.047 | 0.998 | 0.087 | 0.058 | ||||
| 0.005 | 0.865 | 0.004 | 99 | 0.257 | 0.991 | 0.988 | 0.118 | 0.997 | 0.162 | 0.132 | |
| 90 | 0.65 | 0.872 | 0.87 | 0.045 | 0.996 | 0.084 | 0.055 | ||||
| 95 | 0.534 | 0.936 | 0.932 | 0.072 | 0.995 | 0.127 | 0.087 | ||||
| 0.009 | 0.008 | 99 | 0.194 | 0.996 | 0.988 | 0.296 | 0.993 | 0.234 | 0.268 | ||
| 90 | 0.632 | 0.91 | 0.905 | 0.116 | 0.993 | 0.196 | 0.139 | ||||
| 95 | 0.494 | 0.958 | 0.95 | 0.181 | 0.99 | 0.265 | 0.208 | ||||
| Ridge | 0.018 | 0.866 | 0.016 | 99 | 0.219 | 0.994 | 0.98 | 0.404 | 0.986 | 0.284 | 0.346 |
| 90 | 0.647 | 0.902 | 0.901 | 0.03 | 0.998 | 0.057 | 0.037 | ||||
| 95 | 0.517 | 0.952 | 0.95 | 0.047 | 0.998 | 0.087 | 0.058 | ||||
| 0.005 | 0.867 | 0.004 | 99 | 0.257 | 0.991 | 0.988 | 0.117 | 0.997 | 0.161 | 0.132 | |
| 90 | 0.643 | 0.905 | 0.903 | 0.059 | 0.996 | 0.108 | 0.072 | ||||
| 95 | 0.509 | 0.954 | 0.95 | 0.093 | 0.995 | 0.158 | 0.112 | ||||
| 0.009 | 0.008 | 99 | 0.247 | 0.992 | 0.985 | 0.228 | 0.993 | 0.237 | 0.232 | ||
| 90 | 0.635 | 0.91 | 0.905 | 0.117 | 0.993 | 0.197 | 0.139 | ||||
| 95 | 0.494 | 0.958 | 0.95 | 0.181 | 0.99 | 0.265 | 0.207 | ||||
| GLM | 0.018 | 0.868 | 0.016 | 99 | 0.22 | 0.994 | 0.98 | 0.405 | 0.986 | 0.286 | 0.347 |
Abbreviations: Sens: sensitivity; Spec: specificity; Acc: accuracy; GLM: logistic regression; RF: random forests; Ridge: ridge regression.
To investigate the behavior of prediction performance metrics, we varied both (i) the event rate, keeping sample size fixed, and (ii) the absolute number of events, increasing sample size appropriately as the event rate decreases. We varied and considered the three event rates defined above. For each simulation iteration we generated a training data set by sampling covariates for observations using bootstrap sampling with replacement from , and generated outcomes according to each prediction model [30, see, e.g.,]. We then repeated this procedure once for each sample size, event rate, and outcome-generating model to generate an evaluation data set of 1.6 million observations. Performance in this evaluation dataset, a large independent sample, is intended to capture true model performance. For each scenario, defined by event rate, sample size, and algorithm, we conducted 2500 simulation iterations.
2.2.3 Developing clinical prediction models
For each simulation iteration we estimated logistic regression, ridge logistic regression, and random forests prediction models. While logistic regression is a common prediction approach in binary outcome settings, in some rare-event cases, it may be unstable, particularly with a large number of predictors. Ridge regression was used as a penalized logistic regression method [28]. Random forests with probability trees [31] was selected as an example of a more complex prediction modeling approach that estimated interactions between predictors automatically from the data. We did not include any interactions between predictors in models estimated with logistic regression or ridge regression. We used global tuning parameters in model estimation for each simulation iteration because our goal was to understand how the behavior of prediction model performance measures depended on event rate and absolute number of events, not estimate the optimal model for each data set. For information on tuning parameters used for the simulations, see Supplementary Materials Section S1.5.
2.2.4 Assessing clinical prediction model performance
We used 10-fold outcome-stratified cross-validation to estimate performance metrics for each model on the training dataset. At smaller training set sizes and event rates, some folds had no events despite using outcome-stratified sampling (i.e., there were fewer than 10 events in the sampled training data). Folds without events were excluded from calculation of prediction performance. We used the , , and percentiles of predicted probabilities from the training fold for metrics requiring thresholds (all except AUC and Brier score).
Code implementing asymptotic standard errors (ASEs) was not available for many metrics; thus, we used empirical standard errors (ESEs) to describe variability. While the ESE is not available in a data analysis, in simulations it provides a best-case benchmark for variance estimation. The ESE was defined as the standard deviation of estimates across the 2500 simulation replications. For AUC, in addition to the ESE, we calculated ASEs, as code was readily available [32, 33].
We used the evaluation dataset to describe the true prediction performance in our population, computing all performance metrics using the fitted model and thresholds from each replication. The final true population value of each performance metric was defined as the average of the evaluation-set values across the 2500 simulation replications.
We measured reliability for the performance metrics by calculating the empirical bias, variance, and mean squared error (MSE) of the estimated performance metric across the training-set estimated metrics. The evaluation-estimated mean performance metric was treated as the gold standard, i.e., the true performance (see Table 3). We calculated empirical coverage of 95% confidence intervals (CIs) using the ASE for AUC and the ESE for all metrics, comparing the CI from each replication with the evaluation-estimated mean performance metric.
| Performance measure | Formula or measure(s) |
| Empirical bias | |
| Empirical standard error (ESE) | |
| Empirical coverage of 95% CI | |
| Empirical mean squared error |
2.2.5 Computing
All simulations were conducted in R version 4.1.1 [34]. Logistic regression was implemented using the base R stats package, ridge regression using glmnet (version 4.1-4) [35], and random forests using ranger (version 0.14.1) [36]. ASE-based CIs for cross-validated AUC were obtained using cvAUC [33]. The foreach package was used to implement parallel computing [37].
3 Results
We define the effective sample size (ESS) as the amount of information in the entire training sample used to estimate a metric. Therefore, for sensitivity, the ESS is the number of events in the training sample; for specificity, it is the number of non-events. For AUC, the ESS is the minimum of the number of events and non-events; because we consider event rates less than 50%, in our case the ESS is the number of events. The true value of each performance measure depends on the sample size used to estimate the model, event rate, and model-estimation algorithm. The true values for AUC are shown in Table 4; for other estimands see Table LABEL:tab:all_true_values_test.
| Event rate = 0.009 | Event rate = 0.018 | Event rate = 0.005 | |||||||
| n | GLM | RF | Ridge | GLM | RF | Ridge | GLM | RF | Ridge |
| 0.571 | 0.8 | 0.813 | 0.656 | 0.799 | 0.807 | 0.533 | 0.768 | 0.817 | |
| 0.669 | 0.806 | 0.819 | 0.745 | 0.8 | 0.814 | 0.565 | 0.796 | 0.823 | |
| 0.83 | 0.821 | 0.831 | 0.835 | 0.816 | 0.824 | 0.809 | 0.813 | 0.836 | |
| 0.846 | 0.826 | 0.834 | 0.842 | 0.82 | 0.827 | 0.843 | 0.817 | 0.841 | |
| 0.855 | 0.833 | 0.839 | 0.847 | 0.824 | 0.83 | 0.863 | 0.832 | 0.847 | |
Abbreviations: GLM: logistic regression; RF: random forests; Ridge: ridge regression.
Results describing the behavior of AUC estimates are in Figure 1. Bias (top panel) for all algorithms decreases to zero as the ESS increases. For each sample size, bias decreases as the event rate increases, but once a minimum number of events is reached (approximately 1000 events), the event rate no longer impacts bias. The three right farthest dots in each panel show AUC bias of zero at each event rate for a training sample size of 1,000,000. In Figure S1, we show that for each training set size, bias goes to zero for each event rate. These two plots show the ESS, not the event rate, drives AUC bias. This is supported by the AUC influence function form [32]: the estimated AUC (and its asymptotic variance) depends on the number of events and number of non-events, not the event rate itself.
ESE-based AUC 95% CI coverage (bottom panel of Figure 1) is near the nominal level at all sample sizes, while ASE-based CI coverage followed the same pattern as bias, increasing to the nominal level by an ESS of 1000. In Figure S2, we show MSE and CI width for AUC, demonstrating the variance decreases with increasing ESS. We expect poor coverage of ASE-based intervals when the ESS is small, since these are asymptotic intervals [32]. We observed poor AUC behavior (bias and low coverage) for the logistic regression model when the number of predictors exceeded the ESS. This is consistent with work that has shown increased bias and variability in parameter estimates when the number of events per variable is low. A commonly suggested guideline, and one supported by these simulations, is there should be ten or more events per variable to estimate a logistic regression model [38].
Figures 2 and 3 show that the performance of sensitivity and specificity at the 95 percentile of predicted risk are also driven by the ESS (see Supplementary Tables S4–S7 for 90 and 99 percentile, which have a similar pattern). Bias decreases with increasing ESS. ESE-based CI coverage is at the nominal level for sensitivity. For specificity, CI coverage follows a slightly different pattern; while mostly near the nominal level, when the true specificity is close to 1, there is instability in CI coverage estimates.
In contrast to AUC, sensitivity, and specificity, the performance of accuracy, PPV, NPV, , , and Brier score relied in part on the event rate itself, rather than just on the ESS. We display the results for PPV at the 95 percentile of predicted risk in Figure 4; performance for accuracy, NPV, , , and Brier score and the other algorithms is similar to the results for PPV (Supplementary Tables S8–S22, which also include all metrics at the 90 and 99 percentiles). For these estimands, at a given training set size, bias and coverage tended to be worse at the lower event rates.
4 Discussion
Our numerical experiments demonstrate that bias and CI coverage of AUC, sensitivity, and specificity are driven by ESS, not event rate. Stability was reached early for specificity, as we only considered rare events and the ESS for specificity is the number of non-events. For AUC and sensitivity, once a minimum number of events (approximately 1000) was reached, through either a large sample size or higher event rate, bias was near zero and CI coverage was near nominal. Even in the smallest event rate we considered, ( approximately one event per 217 non-events) our simulations show AUC estimates had low bias as long as the training set was sufficiently large (i.e., = 400,000).
In contrast to AUC, the behavior of accuracy, Brier score, PPV, NPV, score, and score all showed some direct reliance on both event rate and training set size. This is not surprising, since the marginal outcome probability is included in the definition of these metrics. However, bias converged to zero and coverage approached the nominal level as the ESS grew, regardless of event rate. While we focused on binary classification in this article, ESS, rather than simply class proportions or number of events, likely plays an important role in bias and CI coverage of prediction performance metrics in settings with multi-class or time-to-event outcomes. Our work also emphasized the need to choose an appropriate prediction modeling approach for the data, which may include dimension reduction. For example, logistic regression did not perform well without a minimum number of events per predictor. Finally, our results highlight the importance of reporting multiple performance metrics with CIs, an approach which has been advocated by others [6, 39, 12, 25].
There are some limitations to our work. First, we did not vary the strength of association between predictors and outcomes. It is possible bias could improve at smaller ESS if there are strong predictors of the outcome. Second, we do not have ASE formulas for metrics besides AUC. While the bootstrap might be tempting to construct standard errors, poor performance of ESE-based intervals, due to the performance metrics being near the boundary values of 0 or 1, suggests the standard bootstrap may not perform well. Bootstrap procedures designed for constructing CIs when estimates are near the boundary are available [40].
Our work shows that AUC estimates are stable and reliable if the ESS is large enough. That is, unbiased AUC estimates are produced if the sample size is large enough even for low event rates. Additionally, AUC ASE-based CIs had nominal coverage at large enough sample sizes, indicating that ASE-based CIs can be reliably used.
Acknowledgments
Research reported in this publication was supported by the National Institute of Mental Health of the National Institutes of Health under award numbers R01 MH125821, U19-MH099201, and U19-MH121738. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
SUPPLEMENTARY MATERIALS
S1 Additional simulation results
S1.1 Tuning parameters
In Table S1, we provide the tuning parameters used in the original outcome regression fits used to generate outcomes in the plasmode simulations presented in the main manuscript.
| Event rate | Algorithm | Tuning parameter(s) |
| GLM | — | |
| GLM | — | |
| GLM | — | |
| Ridge | ||
| Ridge | ||
| Ridge | ||
| RF | mtry = 12, ntree = 50, min.node.size = 10,000 | |
| RF | mtry = 12, ntree = 50, min.node.size = 10,000 | |
| RF | mtry = 12, ntree = 50, min.node.size = 10,000 |
In Table S2, we present the tuning parameters used to fit each algorithm at each sample size and event rate in the simulations.
| Training set size | ||||||
| Algorithm | Event rate | |||||
| Ridge | 0.368710 | 0.112562 | 0.007737 | 0.004101 | 0.000887 | |
| Ridge | 0.271674 | 0.056800 | 0.007044 | 0.001855 | 0.001650 | |
| Ridge | 0.137977 | 0.049776 | 0.003318 | 0.002995 | 0.002987 | |
| Random forests | 1000 | 5000 | 5000 | 5000 | 15000 | |
| Random forests | 1000 | 5000 | 5000 | 5000 | 15000 | |
| Random forests | 1000 | 5000 | 5000 | 5000 | 50000 | |
S1.2 True values of estimands
In Table LABEL:tab:all_true_values_test, we present the true values for each estimand at each combination of sample size, event rate, and algorithm.
| Event rate = 0.009 | Event rate = 0.018 | Event rate = 0.005 | ||||||||
| n | Estimand | GLM | RF | Ridge | GLM | RF | Ridge | GLM | RF | Ridge |
| AUC | 0.571 | 0.8 | 0.813 | 0.656 | 0.799 | 0.807 | 0.533 | 0.768 | 0.817 | |
| AUC | 0.669 | 0.806 | 0.819 | 0.745 | 0.8 | 0.814 | 0.565 | 0.796 | 0.823 | |
| AUC | 0.83 | 0.821 | 0.831 | 0.835 | 0.816 | 0.824 | 0.809 | 0.813 | 0.836 | |
| AUC | 0.846 | 0.826 | 0.834 | 0.842 | 0.82 | 0.827 | 0.843 | 0.817 | 0.841 | |
| AUC | 0.855 | 0.833 | 0.839 | 0.847 | 0.824 | 0.83 | 0.863 | 0.832 | 0.847 | |
| Sensitivity 95% | 0.215 | 0.325 | 0.429 | 0.272 | 0.314 | 0.401 | 0.329 | 0.301 | 0.442 | |
| Sensitivity 95% | 0.309 | 0.336 | 0.438 | 0.35 | 0.319 | 0.411 | 0.229 | 0.325 | 0.453 | |
| Sensitivity 95% | 0.458 | 0.354 | 0.46 | 0.44 | 0.332 | 0.429 | 0.454 | 0.353 | 0.479 | |
| Sensitivity 95% | 0.479 | 0.359 | 0.466 | 0.451 | 0.336 | 0.434 | 0.492 | 0.358 | 0.487 | |
| Sensitivity 95% | 0.495 | 0.369 | 0.475 | 0.46 | 0.34 | 0.44 | 0.522 | 0.372 | 0.5 | |
| Specificity 95% | 0.951 | 0.952 | 0.953 | 0.954 | 0.954 | 0.956 | 0.758 | 0.947 | 0.951 | |
| Specificity 95% | 0.953 | 0.953 | 0.953 | 0.956 | 0.955 | 0.956 | 0.952 | 0.951 | 0.952 | |
| Specificity 95% | 0.954 | 0.953 | 0.954 | 0.957 | 0.955 | 0.957 | 0.952 | 0.951 | 0.952 | |
| Specificity 95% | 0.954 | 0.953 | 0.954 | 0.957 | 0.955 | 0.957 | 0.952 | 0.951 | 0.952 | |
| Specificity 95% | 0.954 | 0.953 | 0.954 | 0.957 | 0.955 | 0.957 | 0.952 | 0.952 | 0.952 | |
| PPV 95% | 0.039 | 0.059 | 0.078 | 0.093 | 0.11 | 0.136 | 0.013 | 0.027 | 0.043 | |
| PPV 95% | 0.057 | 0.061 | 0.08 | 0.12 | 0.112 | 0.14 | 0.023 | 0.03 | 0.044 | |
| PPV 95% | 0.084 | 0.065 | 0.084 | 0.15 | 0.117 | 0.146 | 0.044 | 0.033 | 0.046 | |
| PPV 95% | 0.088 | 0.066 | 0.085 | 0.153 | 0.118 | 0.148 | 0.048 | 0.034 | 0.047 | |
| PPV 95% | 0.091 | 0.068 | 0.087 | 0.156 | 0.119 | 0.15 | 0.051 | 0.035 | 0.048 | |
| Sensitivity 90% | 0.29 | 0.47 | 0.539 | 0.368 | 0.458 | 0.516 | 0.409 | 0.438 | 0.552 | |
| Sensitivity 90% | 0.405 | 0.481 | 0.548 | 0.463 | 0.461 | 0.527 | 0.305 | 0.469 | 0.561 | |
| Sensitivity 90% | 0.575 | 0.502 | 0.572 | 0.564 | 0.478 | 0.547 | 0.565 | 0.5 | 0.587 | |
| Sensitivity 90% | 0.596 | 0.511 | 0.578 | 0.575 | 0.489 | 0.553 | 0.605 | 0.507 | 0.596 | |
| Sensitivity 90% | 0.613 | 0.524 | 0.587 | 0.584 | 0.496 | 0.558 | 0.636 | 0.525 | 0.608 | |
| Sensitivity 99% | 0.113 | 0.128 | 0.221 | 0.132 | 0.118 | 0.189 | 0.103 | 0.107 | 0.234 | |
| Sensitivity 99% | 0.156 | 0.14 | 0.231 | 0.165 | 0.126 | 0.198 | 0.117 | 0.129 | 0.249 | |
| Sensitivity 99% | 0.242 | 0.147 | 0.249 | 0.213 | 0.13 | 0.211 | 0.248 | 0.148 | 0.271 | |
| Sensitivity 99% | 0.257 | 0.151 | 0.254 | 0.221 | 0.133 | 0.215 | 0.276 | 0.15 | 0.279 | |
| Sensitivity 99% | 0.271 | 0.152 | 0.261 | 0.228 | 0.133 | 0.22 | 0.302 | 0.155 | 0.291 | |
| Specificity 90% | 0.902 | 0.903 | 0.903 | 0.906 | 0.906 | 0.907 | 0.679 | 0.896 | 0.902 | |
| Specificity 90% | 0.904 | 0.903 | 0.904 | 0.907 | 0.906 | 0.907 | 0.902 | 0.901 | 0.902 | |
| Specificity 90% | 0.904 | 0.904 | 0.904 | 0.908 | 0.907 | 0.908 | 0.902 | 0.901 | 0.902 | |
| Specificity 90% | 0.905 | 0.904 | 0.904 | 0.908 | 0.907 | 0.908 | 0.902 | 0.901 | 0.902 | |
| Specificity 90% | 0.905 | 0.904 | 0.904 | 0.908 | 0.907 | 0.908 | 0.903 | 0.902 | 0.902 | |
| Specificity 99% | 0.988 | 0.991 | 0.992 | 0.99 | 0.992 | 0.993 | 0.982 | 0.99 | 0.991 | |
| Specificity 99% | 0.991 | 0.991 | 0.992 | 0.992 | 0.992 | 0.993 | 0.99 | 0.99 | 0.991 | |
| Specificity 99% | 0.992 | 0.991 | 0.992 | 0.993 | 0.992 | 0.993 | 0.991 | 0.991 | 0.991 | |
| Specificity 99% | 0.992 | 0.991 | 0.992 | 0.994 | 0.992 | 0.994 | 0.991 | 0.991 | 0.991 | |
| Specificity 99% | 0.992 | 0.991 | 0.992 | 0.994 | 0.992 | 0.994 | 0.991 | 0.991 | 0.991 | |
| PPV 90% | 0.027 | 0.043 | 0.049 | 0.064 | 0.08 | 0.088 | 0.009 | 0.02 | 0.027 | |
| PPV 90% | 0.038 | 0.044 | 0.05 | 0.08 | 0.081 | 0.09 | 0.015 | 0.022 | 0.027 | |
| PPV 90% | 0.053 | 0.046 | 0.052 | 0.096 | 0.084 | 0.093 | 0.028 | 0.023 | 0.028 | |
| PPV 90% | 0.055 | 0.047 | 0.053 | 0.098 | 0.086 | 0.094 | 0.029 | 0.024 | 0.029 | |
| PPV 90% | 0.056 | 0.048 | 0.054 | 0.099 | 0.087 | 0.095 | 0.031 | 0.025 | 0.03 | |
| PPV 99% | 0.081 | 0.118 | 0.204 | 0.192 | 0.21 | 0.327 | 0.028 | 0.054 | 0.114 | |
| PPV 99% | 0.135 | 0.128 | 0.213 | 0.269 | 0.221 | 0.34 | 0.053 | 0.061 | 0.121 | |
| PPV 99% | 0.221 | 0.136 | 0.228 | 0.362 | 0.229 | 0.36 | 0.121 | 0.07 | 0.132 | |
| PPV 99% | 0.235 | 0.138 | 0.232 | 0.375 | 0.233 | 0.367 | 0.134 | 0.071 | 0.136 | |
| PPV 99% | 0.248 | 0.14 | 0.239 | 0.387 | 0.234 | 0.375 | 0.147 | 0.073 | 0.141 | |
| F0.5 90% | 0.033 | 0.052 | 0.06 | 0.076 | 0.096 | 0.105 | 0.011 | 0.024 | 0.033 | |
| F0.5 90% | 0.046 | 0.054 | 0.061 | 0.095 | 0.097 | 0.108 | 0.019 | 0.027 | 0.034 | |
| F0.5 90% | 0.064 | 0.056 | 0.064 | 0.115 | 0.1 | 0.112 | 0.034 | 0.029 | 0.035 | |
| F0.5 90% | 0.067 | 0.057 | 0.065 | 0.117 | 0.103 | 0.113 | 0.036 | 0.029 | 0.036 | |
| F0.5 90% | 0.069 | 0.059 | 0.066 | 0.119 | 0.104 | 0.114 | 0.038 | 0.031 | 0.036 | |
| F0.5 95% | 0.047 | 0.071 | 0.093 | 0.107 | 0.126 | 0.156 | 0.016 | 0.033 | 0.052 | |
| F0.5 95% | 0.069 | 0.073 | 0.096 | 0.139 | 0.129 | 0.161 | 0.028 | 0.037 | 0.054 | |
| F0.5 95% | 0.1 | 0.078 | 0.101 | 0.172 | 0.134 | 0.168 | 0.054 | 0.04 | 0.057 | |
| F0.5 95% | 0.105 | 0.079 | 0.102 | 0.177 | 0.136 | 0.171 | 0.058 | 0.041 | 0.058 | |
| F0.5 95% | 0.108 | 0.081 | 0.104 | 0.18 | 0.137 | 0.173 | 0.062 | 0.043 | 0.059 | |
| F0.5 99% | 0.085 | 0.12 | 0.207 | 0.175 | 0.181 | 0.284 | 0.033 | 0.06 | 0.127 | |
| F0.5 99% | 0.139 | 0.13 | 0.216 | 0.238 | 0.192 | 0.296 | 0.06 | 0.068 | 0.135 | |
| F0.5 99% | 0.225 | 0.138 | 0.232 | 0.318 | 0.199 | 0.315 | 0.135 | 0.078 | 0.147 | |
| F0.5 99% | 0.239 | 0.141 | 0.236 | 0.329 | 0.202 | 0.322 | 0.149 | 0.079 | 0.151 | |
| F0.5 99% | 0.252 | 0.142 | 0.243 | 0.34 | 0.203 | 0.328 | 0.164 | 0.082 | 0.157 | |
| F1 90% | 0.049 | 0.078 | 0.09 | 0.109 | 0.136 | 0.15 | 0.018 | 0.038 | 0.051 | |
| F1 90% | 0.069 | 0.081 | 0.092 | 0.136 | 0.137 | 0.153 | 0.029 | 0.042 | 0.052 | |
| F1 90% | 0.097 | 0.084 | 0.096 | 0.164 | 0.143 | 0.159 | 0.052 | 0.045 | 0.054 | |
| F1 90% | 0.1 | 0.086 | 0.097 | 0.167 | 0.146 | 0.161 | 0.056 | 0.045 | 0.055 | |
| F1 90% | 0.103 | 0.088 | 0.098 | 0.17 | 0.148 | 0.163 | 0.059 | 0.047 | 0.056 | |
| F1 95% | 0.066 | 0.1 | 0.132 | 0.139 | 0.162 | 0.203 | 0.024 | 0.05 | 0.078 | |
| F1 95% | 0.097 | 0.104 | 0.135 | 0.179 | 0.166 | 0.209 | 0.041 | 0.056 | 0.08 | |
| F1 95% | 0.142 | 0.11 | 0.142 | 0.223 | 0.173 | 0.218 | 0.081 | 0.061 | 0.085 | |
| F1 95% | 0.148 | 0.111 | 0.144 | 0.229 | 0.175 | 0.221 | 0.087 | 0.061 | 0.086 | |
| F1 95% | 0.153 | 0.114 | 0.147 | 0.233 | 0.177 | 0.224 | 0.092 | 0.064 | 0.088 | |
| F1 99% | 0.094 | 0.123 | 0.211 | 0.156 | 0.15 | 0.239 | 0.044 | 0.071 | 0.153 | |
| F1 99% | 0.144 | 0.133 | 0.221 | 0.204 | 0.16 | 0.249 | 0.073 | 0.082 | 0.163 | |
| F1 99% | 0.231 | 0.141 | 0.238 | 0.268 | 0.166 | 0.266 | 0.162 | 0.095 | 0.177 | |
| F1 99% | 0.245 | 0.144 | 0.242 | 0.278 | 0.169 | 0.272 | 0.18 | 0.096 | 0.182 | |
| F1 99% | 0.259 | 0.146 | 0.25 | 0.287 | 0.17 | 0.277 | 0.198 | 0.099 | 0.19 | |
| NPV 90% | 0.993 | 0.995 | 0.995 | 0.988 | 0.989 | 0.991 | 0.996 | 0.997 | 0.998 | |
| NPV 90% | 0.994 | 0.995 | 0.995 | 0.99 | 0.989 | 0.991 | 0.996 | 0.997 | 0.998 | |
| NPV 90% | 0.996 | 0.995 | 0.996 | 0.992 | 0.99 | 0.991 | 0.998 | 0.997 | 0.998 | |
| NPV 90% | 0.996 | 0.995 | 0.996 | 0.992 | 0.99 | 0.992 | 0.998 | 0.997 | 0.998 | |
| NPV 90% | 0.996 | 0.995 | 0.996 | 0.992 | 0.99 | 0.992 | 0.998 | 0.998 | 0.998 | |
| NPV 95% | 0.992 | 0.993 | 0.994 | 0.987 | 0.987 | 0.989 | 0.996 | 0.997 | 0.997 | |
| NPV 95% | 0.993 | 0.994 | 0.995 | 0.988 | 0.987 | 0.989 | 0.996 | 0.997 | 0.997 | |
| NPV 95% | 0.995 | 0.994 | 0.995 | 0.99 | 0.988 | 0.99 | 0.997 | 0.997 | 0.997 | |
| NPV 95% | 0.995 | 0.994 | 0.995 | 0.99 | 0.988 | 0.99 | 0.997 | 0.997 | 0.997 | |
| NPV 95% | 0.995 | 0.994 | 0.995 | 0.99 | 0.988 | 0.99 | 0.998 | 0.997 | 0.997 | |
| NPV 99% | 0.992 | 0.992 | 0.993 | 0.985 | 0.984 | 0.986 | 0.996 | 0.996 | 0.996 | |
| NPV 99% | 0.992 | 0.992 | 0.993 | 0.986 | 0.984 | 0.986 | 0.996 | 0.996 | 0.996 | |
| NPV 99% | 0.993 | 0.992 | 0.993 | 0.986 | 0.985 | 0.986 | 0.996 | 0.996 | 0.996 | |
| NPV 99% | 0.993 | 0.992 | 0.993 | 0.987 | 0.985 | 0.986 | 0.996 | 0.996 | 0.996 | |
| NPV 99% | 0.993 | 0.992 | 0.993 | 0.987 | 0.985 | 0.987 | 0.997 | 0.996 | 0.997 | |
| Accuracy 90% | 0.897 | 0.899 | 0.9 | 0.897 | 0.898 | 0.9 | 0.677 | 0.894 | 0.9 | |
| Accuracy 90% | 0.899 | 0.899 | 0.901 | 0.9 | 0.898 | 0.901 | 0.9 | 0.899 | 0.9 | |
| Accuracy 90% | 0.901 | 0.9 | 0.901 | 0.902 | 0.899 | 0.902 | 0.901 | 0.9 | 0.901 | |
| Accuracy 90% | 0.902 | 0.9 | 0.901 | 0.903 | 0.9 | 0.902 | 0.901 | 0.9 | 0.901 | |
| Accuracy 90% | 0.902 | 0.9 | 0.902 | 0.903 | 0.9 | 0.902 | 0.901 | 0.9 | 0.901 | |
| Accuracy 95% | 0.944 | 0.946 | 0.948 | 0.943 | 0.943 | 0.946 | 0.756 | 0.944 | 0.949 | |
| Accuracy 95% | 0.947 | 0.947 | 0.949 | 0.945 | 0.944 | 0.947 | 0.948 | 0.948 | 0.949 | |
| Accuracy 95% | 0.949 | 0.947 | 0.949 | 0.948 | 0.944 | 0.948 | 0.95 | 0.948 | 0.95 | |
| Accuracy 95% | 0.95 | 0.947 | 0.949 | 0.948 | 0.944 | 0.948 | 0.95 | 0.948 | 0.95 | |
| Accuracy 95% | 0.95 | 0.948 | 0.95 | 0.949 | 0.944 | 0.948 | 0.95 | 0.949 | 0.95 | |
| Accuracy 99% | 0.98 | 0.983 | 0.985 | 0.976 | 0.977 | 0.979 | 0.978 | 0.986 | 0.987 | |
| Accuracy 99% | 0.983 | 0.983 | 0.985 | 0.978 | 0.977 | 0.98 | 0.986 | 0.986 | 0.988 | |
| Accuracy 99% | 0.985 | 0.984 | 0.985 | 0.98 | 0.977 | 0.98 | 0.988 | 0.987 | 0.988 | |
| Accuracy 99% | 0.986 | 0.984 | 0.985 | 0.98 | 0.977 | 0.98 | 0.988 | 0.987 | 0.988 | |
| Accuracy 99% | 0.986 | 0.984 | 0.986 | 0.981 | 0.977 | 0.98 | 0.988 | 0.987 | 0.988 | |
| Brier score | 0.013 | 0.009 | 0.009 | 0.019 | 0.017 | 0.015 | 0.015 | 0.009 | 0.005 | |
| Brier score | 0.01 | 0.009 | 0.008 | 0.017 | 0.017 | 0.015 | 0.006 | 0.011 | 0.005 | |
| Brier score | 0.008 | 0.009 | 0.008 | 0.015 | 0.016 | 0.015 | 0.005 | 0.01 | 0.005 | |
| Brier score | 0.008 | 0.009 | 0.008 | 0.015 | 0.016 | 0.015 | 0.005 | 0.01 | 0.005 | |
| Brier score | 0.008 | 0.009 | 0.008 | 0.015 | 0.016 | 0.015 | 0.004 | 0.005 | 0.004 | |
| • Abbreviations: GLM: logistic regression; RF: random forests; Ridge: ridge regression. | ||||||||||
S1.3 Behavior of AUC
In Figure S1, we show bias and coverage for estimating AUC versus training set size. The rows are the event rates, while the columns are the algorithms. We see that the trends in bias and coverage do not depend on event rate, but instead depend on the number of events.
In Figure S2, we show the empirical mean squared error and confidence interval width for estimating AUC.
S1.4 Behavior of other prediction metrics
We present behavior for sensitivity, specificity, and PPV at the 90th and 99th percentiles of predicted risk; accuracy, NPV, F1, and F0.5 at the 90th, 95th, and 99th percentiles of predicted risk; and the Brier score in Tables S4–S22.
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.045 | 0.065 | 0.947 | 1.131 | ||
| GLM | 46 | -0.02 | 0.007 | 0.940 | 0.307 | ||
| GLM | 230 | -0.008 | 0.001 | 0.944 | 0.136 | ||
| GLM | 460 | -0.004 | 0.001 | 0.948 | 0.090 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.944 | 0.028 | ||
| GLM | 46 | -0.013 | 0.008 | 0.959 | 0.351 | ||
| GLM | 92 | -0.019 | 0.004 | 0.938 | 0.235 | ||
| GLM | 460 | -0.005 | 0.001 | 0.945 | 0.099 | ||
| GLM | 919 | -0.001 | 0.001 | 0.952 | 0.065 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.954 | 0.020 | ||
| GLM | 92 | -0.014 | 0.004 | 0.947 | 0.243 | ||
| GLM | 184 | -0.011 | 0.002 | 0.946 | 0.172 | ||
| GLM | 919 | -0.002 | 0.001 | 0.941 | 0.067 | ||
| GLM | 1837 | -0.001 | 0.001 | 0.952 | 0.047 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.945 | 0.015 | ||
| RF | 23 | -0.006 | 0.014 | 0.951 | 0.487 | ||
| RF | 46 | 0.001 | 0.008 | 0.947 | 0.319 | ||
| RF | 230 | 0.002 | 0.002 | 0.950 | 0.134 | ||
| RF | 460 | 0.002 | 0.002 | 0.948 | 0.097 | ||
| RF | 4591 | 0.001 | 0.001 | 0.946 | 0.028 | ||
| RF | 46 | -0.002 | 0.006 | 0.952 | 0.317 | ||
| RF | 92 | -0.002 | 0.003 | 0.950 | 0.210 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.089 | ||
| RF | 919 | -0.001 | 0.001 | 0.952 | 0.064 | ||
| RF | 9181 | 0.001 | 0.001 | 0.944 | 0.019 | ||
| RF | 92 | -0.002 | 0.003 | 0.953 | 0.216 | ||
| RF | 184 | 0.001 | 0.002 | 0.949 | 0.149 | ||
| RF | 919 | 0.001 | 0.001 | 0.950 | 0.063 | ||
| RF | 1837 | -0.002 | 0.001 | 0.939 | 0.045 | ||
| RF | 18361 | 0.001 | 0.001 | 0.952 | 0.015 | ||
| Ridge | 23 | 0.001 | 0.012 | 0.950 | 0.426 | ||
| Ridge | 46 | -0.002 | 0.005 | 0.946 | 0.287 | ||
| Ridge | 230 | -0.003 | 0.001 | 0.952 | 0.128 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.088 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.952 | 0.028 | ||
| Ridge | 46 | 0.001 | 0.006 | 0.947 | 0.301 | ||
| Ridge | 92 | -0.001 | 0.003 | 0.952 | 0.206 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.950 | 0.091 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.063 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.952 | 0.020 | ||
| Ridge | 92 | -0.003 | 0.003 | 0.947 | 0.214 | ||
| Ridge | 184 | -0.002 | 0.002 | 0.946 | 0.152 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.067 | ||
| Ridge | 1837 | -0.001 | 0.001 | 0.950 | 0.046 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.953 | 0.015 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.009 | 0.005 | 0.958 | 0.301 | ||
| GLM | 46 | -0.008 | 0.003 | 0.952 | 0.193 | ||
| GLM | 230 | -0.006 | 0.001 | 0.946 | 0.116 | ||
| GLM | 460 | -0.003 | 0.001 | 0.949 | 0.080 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.954 | 0.024 | ||
| GLM | 46 | -0.003 | 0.003 | 0.954 | 0.205 | ||
| GLM | 92 | -0.007 | 0.002 | 0.949 | 0.161 | ||
| GLM | 460 | -0.003 | 0.001 | 0.945 | 0.080 | ||
| GLM | 919 | -0.001 | 0.001 | 0.945 | 0.054 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.017 | ||
| GLM | 92 | -0.004 | 0.002 | 0.954 | 0.154 | ||
| GLM | 184 | -0.004 | 0.001 | 0.950 | 0.117 | ||
| GLM | 919 | -0.002 | 0.001 | 0.946 | 0.050 | ||
| GLM | 1837 | -0.001 | 0.001 | 0.946 | 0.034 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.951 | 0.010 | ||
| RF | 23 | -0.001 | 0.005 | 0.958 | 0.317 | ||
| RF | 46 | 0.001 | 0.003 | 0.966 | 0.234 | ||
| RF | 230 | 0.001 | 0.001 | 0.957 | 0.097 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.068 | ||
| RF | 4591 | 0.001 | 0.001 | 0.949 | 0.020 | ||
| RF | 46 | -0.002 | 0.003 | 0.960 | 0.232 | ||
| RF | 92 | -0.002 | 0.002 | 0.948 | 0.160 | ||
| RF | 460 | 0.001 | 0.001 | 0.946 | 0.063 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.044 | ||
| RF | 9181 | 0.001 | 0.001 | 0.945 | 0.014 | ||
| RF | 92 | 0.001 | 0.001 | 0.956 | 0.148 | ||
| RF | 184 | -0.001 | 0.001 | 0.950 | 0.102 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.041 | ||
| RF | 1837 | 0.001 | 0.001 | 0.953 | 0.028 | ||
| RF | 18361 | 0.001 | 0.001 | 0.952 | 0.009 | ||
| Ridge | 23 | -0.004 | 0.008 | 0.962 | 0.369 | ||
| Ridge | 46 | -0.003 | 0.004 | 0.952 | 0.257 | ||
| Ridge | 230 | -0.002 | 0.001 | 0.951 | 0.112 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.948 | 0.080 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.950 | 0.025 | ||
| Ridge | 46 | -0.001 | 0.004 | 0.953 | 0.244 | ||
| Ridge | 92 | -0.001 | 0.002 | 0.952 | 0.173 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.076 | ||
| Ridge | 919 | -0.001 | 0.001 | 0.951 | 0.054 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.952 | 0.016 | ||
| Ridge | 92 | -0.003 | 0.002 | 0.952 | 0.159 | ||
| Ridge | 184 | -0.001 | 0.001 | 0.955 | 0.113 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.953 | 0.049 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.947 | 0.034 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.952 | 0.011 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.048 | 0.075 | 0.952 | 1.215 | ||
| GLM | 46 | 0.001 | 0.001 | 0.998 | 0.034 | ||
| GLM | 230 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| GLM | 460 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.950 | 0.000 | ||
| GLM | 46 | -0.004 | 0.001 | 0.980 | 0.166 | ||
| GLM | 92 | 0.001 | 0.001 | 0.949 | 0.007 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| GLM | 919 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.928 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.947 | 0.011 | ||
| GLM | 184 | 0.001 | 0.001 | 0.949 | 0.007 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.935 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.937 | 0.011 | ||
| RF | 46 | 0.001 | 0.001 | 0.949 | 0.006 | ||
| RF | 230 | 0.001 | 0.001 | 0.953 | 0.002 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 4591 | 0.001 | 0.001 | 0.946 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.946 | 0.008 | ||
| RF | 92 | 0.001 | 0.001 | 0.948 | 0.005 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.955 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.952 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.009 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.006 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.945 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.945 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.946 | 0.006 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.939 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.948 | 0.006 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.944 | 0.005 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.954 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.944 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.952 | 0.007 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.952 | 0.005 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.944 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.927 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.003 | 0.001 | 0.900 | 0.019 | ||
| GLM | 46 | 0.001 | 0.001 | 0.932 | 0.003 | ||
| GLM | 230 | 0.001 | 0.001 | 0.940 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.917 | 0.000 | ||
| GLM | 46 | -0.001 | 0.001 | 0.936 | 0.007 | ||
| GLM | 92 | 0.001 | 0.001 | 0.944 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.938 | 0.006 | ||
| GLM | 184 | 0.001 | 0.001 | 0.943 | 0.003 | ||
| GLM | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.897 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.786 | 0.004 | ||
| RF | 46 | 0.001 | 0.001 | 0.942 | 0.002 | ||
| RF | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.944 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.935 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.950 | 0.004 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.002 | ||
| RF | 460 | 0.001 | 0.001 | 0.954 | 0.001 | ||
| RF | 919 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.925 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.003 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 1837 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| RF | 18361 | 0.001 | 0.001 | 0.961 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.942 | 0.003 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.955 | 0.002 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.861 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.944 | 0.003 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.958 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.941 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.922 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.942 | 0.003 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.941 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.888 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.048 | 0.074 | 0.952 | 1.204 | ||
| GLM | 46 | 0.001 | 0.001 | 0.998 | 0.033 | ||
| GLM | 230 | 0.001 | 0.001 | 0.950 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.943 | 0.000 | ||
| GLM | 46 | -0.004 | 0.001 | 0.981 | 0.163 | ||
| GLM | 92 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| GLM | 460 | 0.001 | 0.001 | 0.946 | 0.003 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.930 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.953 | 0.012 | ||
| GLM | 184 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.003 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.943 | 0.001 | ||
| RF | 23 | 0.001 | 0.001 | 0.937 | 0.011 | ||
| RF | 46 | 0.001 | 0.001 | 0.954 | 0.006 | ||
| RF | 230 | 0.001 | 0.001 | 0.948 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.944 | 0.002 | ||
| RF | 4591 | 0.001 | 0.001 | 0.955 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.951 | 0.009 | ||
| RF | 92 | 0.001 | 0.001 | 0.956 | 0.006 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| RF | 9181 | 0.001 | 0.001 | 0.906 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.952 | 0.010 | ||
| RF | 184 | 0.001 | 0.001 | 0.951 | 0.007 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.003 | ||
| RF | 1837 | 0.001 | 0.001 | 0.941 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.945 | 0.001 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.952 | 0.007 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.948 | 0.005 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.917 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.945 | 0.007 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.947 | 0.005 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.953 | 0.002 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.957 | 0.009 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.950 | 0.006 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.946 | 0.003 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.947 | 0.001 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.039 | 0.064 | 0.917 | 1.188 | ||
| GLM | 46 | 0.001 | 0.001 | 0.999 | 0.028 | ||
| GLM | 230 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| GLM | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.945 | 0.000 | ||
| GLM | 46 | -0.003 | 0.001 | 0.983 | 0.142 | ||
| GLM | 92 | 0.001 | 0.001 | 0.951 | 0.006 | ||
| GLM | 460 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| GLM | 919 | 0.001 | 0.001 | 0.942 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.956 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.941 | 0.009 | ||
| GLM | 184 | 0.001 | 0.001 | 0.944 | 0.006 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.941 | 0.001 | ||
| RF | 23 | 0.001 | 0.001 | 0.923 | 0.008 | ||
| RF | 46 | 0.001 | 0.001 | 0.954 | 0.005 | ||
| RF | 230 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| RF | 460 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.944 | 0.008 | ||
| RF | 92 | 0.001 | 0.001 | 0.951 | 0.005 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.947 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.946 | 0.009 | ||
| RF | 184 | 0.001 | 0.001 | 0.946 | 0.006 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.003 | ||
| RF | 1837 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.955 | 0.005 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.946 | 0.004 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.947 | 0.006 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.948 | 0.005 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.943 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.945 | 0.006 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.945 | 0.002 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.936 | 0.001 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.003 | 0.001 | 0.890 | 0.017 | ||
| GLM | 46 | 0.001 | 0.001 | 0.930 | 0.004 | ||
| GLM | 230 | 0.001 | 0.001 | 0.938 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.954 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.941 | 0.000 | ||
| GLM | 46 | -0.001 | 0.001 | 0.921 | 0.006 | ||
| GLM | 92 | 0.001 | 0.001 | 0.942 | 0.004 | ||
| GLM | 460 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| GLM | 919 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.938 | 0.008 | ||
| GLM | 184 | 0.001 | 0.001 | 0.952 | 0.005 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.925 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.849 | 0.005 | ||
| RF | 46 | 0.001 | 0.001 | 0.961 | 0.004 | ||
| RF | 230 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.942 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.887 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.953 | 0.006 | ||
| RF | 92 | 0.001 | 0.001 | 0.954 | 0.004 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.943 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.949 | 0.008 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.005 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.949 | 0.004 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.958 | 0.003 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.922 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.956 | 0.005 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.947 | 0.004 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.945 | 0.002 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.933 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.956 | 0.007 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.946 | 0.005 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.956 | 0.002 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.930 | 0.001 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.950 | 0.021 | ||
| GLM | 46 | 0.001 | 0.001 | 0.950 | 0.020 | ||
| GLM | 230 | 0.001 | 0.001 | 0.951 | 0.010 | ||
| GLM | 460 | 0.001 | 0.001 | 0.951 | 0.007 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| GLM | 46 | -0.001 | 0.001 | 0.952 | 0.037 | ||
| GLM | 92 | -0.001 | 0.001 | 0.945 | 0.029 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.014 | ||
| GLM | 919 | 0.001 | 0.001 | 0.952 | 0.009 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.952 | 0.003 | ||
| GLM | 92 | -0.002 | 0.001 | 0.948 | 0.055 | ||
| GLM | 184 | -0.001 | 0.001 | 0.946 | 0.041 | ||
| GLM | 919 | 0.001 | 0.001 | 0.948 | 0.017 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.012 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.953 | 0.004 | ||
| RF | 23 | 0.001 | 0.001 | 0.954 | 0.028 | ||
| RF | 46 | 0.001 | 0.001 | 0.942 | 0.020 | ||
| RF | 230 | 0.001 | 0.001 | 0.952 | 0.009 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.006 | ||
| RF | 4591 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 46 | 0.001 | 0.001 | 0.955 | 0.039 | ||
| RF | 92 | 0.001 | 0.001 | 0.951 | 0.027 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.012 | ||
| RF | 919 | 0.001 | 0.001 | 0.942 | 0.008 | ||
| RF | 9181 | 0.001 | 0.001 | 0.943 | 0.003 | ||
| RF | 92 | 0.001 | 0.001 | 0.954 | 0.051 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.035 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.015 | ||
| RF | 1837 | 0.001 | 0.001 | 0.951 | 0.011 | ||
| RF | 18361 | 0.001 | 0.001 | 0.951 | 0.004 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.949 | 0.029 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.952 | 0.021 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.944 | 0.010 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.007 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.955 | 0.002 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.955 | 0.040 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.955 | 0.029 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.954 | 0.013 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.951 | 0.009 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.948 | 0.003 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.948 | 0.052 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.948 | 0.037 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.017 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.947 | 0.011 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.952 | 0.004 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.001 | 0.001 | 0.960 | 0.079 | ||
| GLM | 46 | -0.004 | 0.001 | 0.959 | 0.102 | ||
| GLM | 230 | -0.002 | 0.001 | 0.952 | 0.066 | ||
| GLM | 460 | 0.001 | 0.001 | 0.952 | 0.046 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.948 | 0.014 | ||
| GLM | 46 | -0.005 | 0.002 | 0.964 | 0.167 | ||
| GLM | 92 | -0.006 | 0.001 | 0.951 | 0.164 | ||
| GLM | 460 | -0.002 | 0.001 | 0.944 | 0.085 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.057 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.018 | ||
| GLM | 92 | -0.008 | 0.004 | 0.963 | 0.263 | ||
| GLM | 184 | -0.006 | 0.003 | 0.948 | 0.216 | ||
| GLM | 919 | -0.003 | 0.001 | 0.950 | 0.096 | ||
| GLM | 1837 | -0.001 | 0.001 | 0.952 | 0.066 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.951 | 0.020 | ||
| RF | 23 | 0.002 | 0.002 | 0.954 | 0.187 | ||
| RF | 46 | 0.001 | 0.001 | 0.965 | 0.123 | ||
| RF | 230 | 0.001 | 0.001 | 0.958 | 0.050 | ||
| RF | 460 | 0.001 | 0.001 | 0.947 | 0.035 | ||
| RF | 4591 | 0.001 | 0.001 | 0.940 | 0.010 | ||
| RF | 46 | 0.002 | 0.003 | 0.958 | 0.252 | ||
| RF | 92 | 0.001 | 0.001 | 0.950 | 0.165 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.063 | ||
| RF | 919 | 0.001 | 0.001 | 0.951 | 0.044 | ||
| RF | 9181 | 0.001 | 0.001 | 0.948 | 0.014 | ||
| RF | 92 | 0.002 | 0.005 | 0.956 | 0.302 | ||
| RF | 184 | 0.001 | 0.002 | 0.947 | 0.199 | ||
| RF | 919 | 0.001 | 0.001 | 0.954 | 0.076 | ||
| RF | 1837 | 0.001 | 0.001 | 0.956 | 0.053 | ||
| RF | 18361 | 0.001 | 0.001 | 0.958 | 0.017 | ||
| Ridge | 23 | 0.002 | 0.003 | 0.948 | 0.224 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.948 | 0.152 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.950 | 0.065 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.946 | 0.046 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.944 | 0.014 | ||
| Ridge | 46 | 0.005 | 0.005 | 0.948 | 0.285 | ||
| Ridge | 92 | 0.004 | 0.002 | 0.946 | 0.193 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.950 | 0.081 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.951 | 0.058 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.948 | 0.017 | ||
| Ridge | 92 | 0.002 | 0.006 | 0.954 | 0.332 | ||
| Ridge | 184 | 0.004 | 0.003 | 0.957 | 0.219 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.951 | 0.092 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.946 | 0.064 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.952 | 0.020 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.958 | 0.004 | ||
| GLM | 46 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| GLM | 230 | 0.001 | 0.001 | 0.928 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.932 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.892 | 0.000 | ||
| GLM | 46 | 0.001 | 0.001 | 0.939 | 0.005 | ||
| GLM | 92 | 0.001 | 0.001 | 0.942 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.944 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.911 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.949 | 0.006 | ||
| GLM | 184 | 0.001 | 0.001 | 0.945 | 0.004 | ||
| GLM | 919 | 0.001 | 0.001 | 0.942 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.917 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.939 | 0.003 | ||
| RF | 46 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| RF | 230 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.937 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.952 | 0.004 | ||
| RF | 92 | 0.001 | 0.001 | 0.947 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| RF | 919 | 0.001 | 0.001 | 0.952 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.885 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.006 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.004 | ||
| RF | 919 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.955 | 0.001 | ||
| RF | 18361 | 0.001 | 0.001 | 0.936 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.956 | 0.003 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.943 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.943 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.956 | 0.004 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.947 | 0.003 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.947 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.956 | 0.006 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.949 | 0.004 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.955 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.939 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.937 | 0.004 | ||
| GLM | 46 | 0.001 | 0.001 | 0.946 | 0.002 | ||
| GLM | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.943 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.858 | 0.000 | ||
| GLM | 46 | 0.001 | 0.001 | 0.947 | 0.005 | ||
| GLM | 92 | 0.001 | 0.001 | 0.944 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.942 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.920 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.946 | 0.006 | ||
| GLM | 184 | 0.001 | 0.001 | 0.945 | 0.004 | ||
| GLM | 919 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.922 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.947 | 0.003 | ||
| RF | 46 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| RF | 230 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.939 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.944 | 0.005 | ||
| RF | 92 | 0.001 | 0.001 | 0.949 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.944 | 0.001 | ||
| RF | 919 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.937 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.957 | 0.007 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.005 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| RF | 18361 | 0.001 | 0.001 | 0.956 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.945 | 0.003 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.947 | 0.002 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.941 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.856 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.949 | 0.004 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.947 | 0.003 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.942 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.926 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.950 | 0.006 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.946 | 0.004 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.917 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.950 | 0.004 | ||
| GLM | 46 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| GLM | 230 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.928 | 0.000 | ||
| GLM | 46 | 0.001 | 0.001 | 0.954 | 0.005 | ||
| GLM | 92 | 0.001 | 0.001 | 0.956 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.951 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.954 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.926 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.946 | 0.007 | ||
| GLM | 184 | 0.001 | 0.001 | 0.941 | 0.004 | ||
| GLM | 919 | 0.001 | 0.001 | 0.954 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.930 | 0.000 | ||
| RF | 23 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| RF | 46 | 0.001 | 0.001 | 0.955 | 0.003 | ||
| RF | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.938 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.950 | 0.005 | ||
| RF | 92 | 0.001 | 0.001 | 0.954 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.924 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.951 | 0.007 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.005 | ||
| RF | 919 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.950 | 0.002 | ||
| RF | 18361 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.958 | 0.003 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.937 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.960 | 0.005 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.953 | 0.003 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.937 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.954 | 0.007 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.946 | 0.005 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.953 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.943 | 0.000 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.950 | 0.039 | ||
| GLM | 46 | -0.002 | 0.001 | 0.944 | 0.037 | ||
| GLM | 230 | 0.001 | 0.001 | 0.951 | 0.019 | ||
| GLM | 460 | 0.001 | 0.001 | 0.951 | 0.013 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| GLM | 46 | -0.002 | 0.001 | 0.951 | 0.065 | ||
| GLM | 92 | -0.003 | 0.001 | 0.946 | 0.052 | ||
| GLM | 460 | 0.001 | 0.001 | 0.949 | 0.025 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.016 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.955 | 0.005 | ||
| GLM | 92 | -0.004 | 0.001 | 0.948 | 0.089 | ||
| GLM | 184 | -0.003 | 0.001 | 0.944 | 0.065 | ||
| GLM | 919 | 0.001 | 0.001 | 0.946 | 0.027 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.019 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.957 | 0.006 | ||
| RF | 23 | 0.001 | 0.001 | 0.952 | 0.052 | ||
| RF | 46 | 0.001 | 0.001 | 0.942 | 0.037 | ||
| RF | 230 | 0.001 | 0.001 | 0.953 | 0.016 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.012 | ||
| RF | 4591 | 0.001 | 0.001 | 0.950 | 0.003 | ||
| RF | 46 | 0.001 | 0.001 | 0.955 | 0.068 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.047 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.020 | ||
| RF | 919 | 0.001 | 0.001 | 0.941 | 0.015 | ||
| RF | 9181 | 0.001 | 0.001 | 0.945 | 0.005 | ||
| RF | 92 | 0.001 | 0.001 | 0.955 | 0.080 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.055 | ||
| RF | 919 | 0.001 | 0.001 | 0.947 | 0.024 | ||
| RF | 1837 | 0.001 | 0.001 | 0.946 | 0.017 | ||
| RF | 18361 | 0.001 | 0.001 | 0.949 | 0.006 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.948 | 0.054 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.952 | 0.039 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.945 | 0.018 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.948 | 0.012 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.954 | 0.004 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.956 | 0.069 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.954 | 0.050 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.953 | 0.023 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.016 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.952 | 0.005 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.946 | 0.082 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.950 | 0.059 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.026 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.948 | 0.018 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.951 | 0.006 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.001 | 0.001 | 0.953 | 0.062 | ||
| GLM | 46 | -0.002 | 0.001 | 0.952 | 0.059 | ||
| GLM | 230 | -0.001 | 0.001 | 0.949 | 0.032 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.022 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.943 | 0.007 | ||
| GLM | 46 | -0.004 | 0.001 | 0.954 | 0.097 | ||
| GLM | 92 | -0.005 | 0.001 | 0.945 | 0.082 | ||
| GLM | 460 | -0.001 | 0.001 | 0.944 | 0.039 | ||
| GLM | 919 | 0.001 | 0.001 | 0.950 | 0.026 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| GLM | 92 | -0.006 | 0.001 | 0.948 | 0.124 | ||
| GLM | 184 | -0.005 | 0.001 | 0.939 | 0.095 | ||
| GLM | 919 | -0.001 | 0.001 | 0.947 | 0.040 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.951 | 0.028 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.946 | 0.009 | ||
| RF | 23 | 0.001 | 0.001 | 0.956 | 0.084 | ||
| RF | 46 | 0.001 | 0.001 | 0.953 | 0.060 | ||
| RF | 230 | 0.001 | 0.001 | 0.945 | 0.026 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.019 | ||
| RF | 4591 | 0.001 | 0.001 | 0.942 | 0.006 | ||
| RF | 46 | 0.001 | 0.001 | 0.952 | 0.107 | ||
| RF | 92 | 0.001 | 0.001 | 0.948 | 0.074 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.031 | ||
| RF | 919 | 0.001 | 0.001 | 0.951 | 0.022 | ||
| RF | 9181 | 0.001 | 0.001 | 0.948 | 0.007 | ||
| RF | 92 | -0.001 | 0.001 | 0.954 | 0.115 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.078 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.034 | ||
| RF | 1837 | 0.001 | 0.001 | 0.945 | 0.024 | ||
| RF | 18361 | 0.001 | 0.001 | 0.941 | 0.008 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.951 | 0.093 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.951 | 0.067 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.952 | 0.030 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.948 | 0.021 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.953 | 0.007 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.954 | 0.112 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.951 | 0.080 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.948 | 0.036 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.947 | 0.025 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.946 | 0.008 | ||
| Ridge | 92 | -0.002 | 0.001 | 0.950 | 0.121 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.955 | 0.088 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.951 | 0.039 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.948 | 0.026 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.950 | 0.009 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | -0.002 | 0.001 | 0.965 | 0.111 | ||
| GLM | 46 | -0.007 | 0.001 | 0.963 | 0.126 | ||
| GLM | 230 | -0.004 | 0.001 | 0.947 | 0.082 | ||
| GLM | 460 | -0.002 | 0.001 | 0.951 | 0.057 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.952 | 0.017 | ||
| GLM | 46 | -0.008 | 0.002 | 0.964 | 0.167 | ||
| GLM | 92 | -0.01 | 0.001 | 0.950 | 0.153 | ||
| GLM | 460 | -0.004 | 0.001 | 0.937 | 0.079 | ||
| GLM | 919 | -0.002 | 0.001 | 0.944 | 0.053 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.951 | 0.017 | ||
| GLM | 92 | -0.012 | 0.002 | 0.950 | 0.177 | ||
| GLM | 184 | -0.009 | 0.001 | 0.940 | 0.143 | ||
| GLM | 919 | -0.004 | 0.001 | 0.943 | 0.063 | ||
| GLM | 1837 | -0.002 | 0.001 | 0.950 | 0.043 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.956 | 0.013 | ||
| RF | 23 | -0.003 | 0.002 | 0.965 | 0.209 | ||
| RF | 46 | -0.002 | 0.001 | 0.965 | 0.152 | ||
| RF | 230 | 0.001 | 0.001 | 0.959 | 0.064 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.045 | ||
| RF | 4591 | 0.001 | 0.001 | 0.943 | 0.013 | ||
| RF | 46 | -0.006 | 0.003 | 0.962 | 0.219 | ||
| RF | 92 | -0.005 | 0.001 | 0.947 | 0.153 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.061 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.043 | ||
| RF | 9181 | 0.001 | 0.001 | 0.946 | 0.013 | ||
| RF | 92 | -0.006 | 0.002 | 0.953 | 0.182 | ||
| RF | 184 | -0.005 | 0.001 | 0.950 | 0.127 | ||
| RF | 919 | 0.001 | 0.001 | 0.952 | 0.052 | ||
| RF | 1837 | -0.001 | 0.001 | 0.950 | 0.036 | ||
| RF | 18361 | 0.001 | 0.001 | 0.949 | 0.011 | ||
| Ridge | 23 | -0.008 | 0.004 | 0.947 | 0.247 | ||
| Ridge | 46 | -0.005 | 0.002 | 0.947 | 0.177 | ||
| Ridge | 230 | -0.002 | 0.001 | 0.952 | 0.079 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.943 | 0.056 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.942 | 0.018 | ||
| Ridge | 46 | -0.009 | 0.003 | 0.950 | 0.235 | ||
| Ridge | 92 | -0.005 | 0.002 | 0.944 | 0.169 | ||
| Ridge | 460 | -0.002 | 0.001 | 0.950 | 0.075 | ||
| Ridge | 919 | -0.002 | 0.001 | 0.952 | 0.054 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.955 | 0.016 | ||
| Ridge | 92 | -0.011 | 0.003 | 0.946 | 0.196 | ||
| Ridge | 184 | -0.006 | 0.001 | 0.953 | 0.140 | ||
| Ridge | 919 | -0.002 | 0.001 | 0.954 | 0.061 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.945 | 0.043 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.948 | 0.014 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.009 | 0.001 | 0.860 | 0.041 | ||
| GLM | 46 | 0.003 | 0.001 | 0.894 | 0.020 | ||
| GLM | 230 | 0.001 | 0.001 | 0.951 | 0.013 | ||
| GLM | 460 | 0.001 | 0.001 | 0.952 | 0.009 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.950 | 0.003 | ||
| GLM | 46 | 0.008 | 0.001 | 0.863 | 0.037 | ||
| GLM | 92 | -0.001 | 0.001 | 0.950 | 0.034 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.017 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.011 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.954 | 0.004 | ||
| GLM | 92 | 0.001 | 0.001 | 0.954 | 0.060 | ||
| GLM | 184 | -0.002 | 0.001 | 0.946 | 0.048 | ||
| GLM | 919 | 0.001 | 0.001 | 0.947 | 0.020 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.014 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.955 | 0.004 | ||
| RF | 23 | 0.009 | 0.001 | 0.732 | 0.025 | ||
| RF | 46 | 0.002 | 0.001 | 0.942 | 0.022 | ||
| RF | 230 | 0.001 | 0.001 | 0.953 | 0.011 | ||
| RF | 460 | 0.001 | 0.001 | 0.950 | 0.008 | ||
| RF | 4591 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| RF | 46 | 0.003 | 0.001 | 0.941 | 0.041 | ||
| RF | 92 | 0.001 | 0.001 | 0.952 | 0.032 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.014 | ||
| RF | 919 | 0.001 | 0.001 | 0.942 | 0.010 | ||
| RF | 9181 | 0.001 | 0.001 | 0.943 | 0.003 | ||
| RF | 92 | 0.001 | 0.001 | 0.953 | 0.058 | ||
| RF | 184 | 0.001 | 0.001 | 0.948 | 0.041 | ||
| RF | 919 | 0.001 | 0.001 | 0.946 | 0.018 | ||
| RF | 1837 | 0.001 | 0.001 | 0.950 | 0.013 | ||
| RF | 18361 | 0.001 | 0.001 | 0.948 | 0.004 | ||
| Ridge | 23 | 0.006 | 0.001 | 0.863 | 0.029 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.946 | 0.024 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.944 | 0.012 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.008 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.957 | 0.003 | ||
| Ridge | 46 | 0.002 | 0.001 | 0.942 | 0.045 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.954 | 0.034 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.954 | 0.016 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.011 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.948 | 0.003 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.949 | 0.060 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.948 | 0.044 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.948 | 0.019 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.947 | 0.013 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.951 | 0.004 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.022 | 0.001 | 0.843 | 0.084 | ||
| GLM | 46 | 0.01 | 0.001 | 0.786 | 0.033 | ||
| GLM | 230 | 0.001 | 0.001 | 0.950 | 0.022 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.015 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.945 | 0.005 | ||
| GLM | 46 | 0.022 | 0.001 | 0.730 | 0.061 | ||
| GLM | 92 | 0.001 | 0.001 | 0.951 | 0.055 | ||
| GLM | 460 | 0.001 | 0.001 | 0.944 | 0.029 | ||
| GLM | 919 | 0.001 | 0.001 | 0.952 | 0.020 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.952 | 0.006 | ||
| GLM | 92 | 0.005 | 0.001 | 0.948 | 0.088 | ||
| GLM | 184 | -0.003 | 0.001 | 0.942 | 0.078 | ||
| GLM | 919 | 0.001 | 0.001 | 0.947 | 0.033 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.024 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.947 | 0.007 | ||
| RF | 23 | 0.025 | 0.001 | 0.420 | 0.043 | ||
| RF | 46 | 0.007 | 0.001 | 0.878 | 0.035 | ||
| RF | 230 | 0.001 | 0.001 | 0.947 | 0.018 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.013 | ||
| RF | 4591 | 0.001 | 0.001 | 0.946 | 0.004 | ||
| RF | 46 | 0.015 | 0.001 | 0.863 | 0.065 | ||
| RF | 92 | 0.002 | 0.001 | 0.948 | 0.051 | ||
| RF | 460 | 0.001 | 0.001 | 0.948 | 0.023 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.016 | ||
| RF | 9181 | 0.001 | 0.001 | 0.946 | 0.005 | ||
| RF | 92 | 0.005 | 0.001 | 0.946 | 0.089 | ||
| RF | 184 | 0.001 | 0.001 | 0.952 | 0.064 | ||
| RF | 919 | 0.001 | 0.001 | 0.950 | 0.028 | ||
| RF | 1837 | 0.001 | 0.001 | 0.946 | 0.020 | ||
| RF | 18361 | 0.001 | 0.001 | 0.945 | 0.006 | ||
| Ridge | 23 | 0.018 | 0.001 | 0.747 | 0.051 | ||
| Ridge | 46 | 0.003 | 0.001 | 0.940 | 0.041 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.953 | 0.021 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.949 | 0.015 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.951 | 0.005 | ||
| Ridge | 46 | 0.009 | 0.001 | 0.921 | 0.074 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.950 | 0.058 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.947 | 0.027 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.950 | 0.019 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.945 | 0.006 | ||
| Ridge | 92 | 0.002 | 0.001 | 0.950 | 0.097 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.956 | 0.073 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.954 | 0.032 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.948 | 0.022 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.951 | 0.007 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.096 | 0.012 | 0.606 | 0.195 | ||
| GLM | 46 | 0.067 | 0.005 | 0.431 | 0.116 | ||
| GLM | 230 | -0.003 | 0.001 | 0.947 | 0.071 | ||
| GLM | 460 | -0.001 | 0.001 | 0.953 | 0.050 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.947 | 0.015 | ||
| GLM | 46 | 0.102 | 0.013 | 0.494 | 0.193 | ||
| GLM | 92 | 0.03 | 0.002 | 0.854 | 0.133 | ||
| GLM | 460 | -0.003 | 0.001 | 0.945 | 0.082 | ||
| GLM | 919 | -0.001 | 0.001 | 0.948 | 0.055 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.952 | 0.018 | ||
| GLM | 92 | 0.06 | 0.006 | 0.758 | 0.178 | ||
| GLM | 184 | 0.002 | 0.002 | 0.954 | 0.161 | ||
| GLM | 919 | -0.004 | 0.001 | 0.943 | 0.078 | ||
| GLM | 1837 | -0.002 | 0.001 | 0.950 | 0.053 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.953 | 0.016 | ||
| RF | 23 | 0.186 | 0.041 | 0.364 | 0.309 | ||
| RF | 46 | 0.072 | 0.006 | 0.416 | 0.126 | ||
| RF | 230 | 0.002 | 0.001 | 0.954 | 0.053 | ||
| RF | 460 | 0.001 | 0.001 | 0.946 | 0.038 | ||
| RF | 4591 | 0.001 | 0.001 | 0.942 | 0.011 | ||
| RF | 46 | 0.126 | 0.019 | 0.430 | 0.229 | ||
| RF | 92 | 0.041 | 0.003 | 0.777 | 0.132 | ||
| RF | 460 | 0.001 | 0.001 | 0.951 | 0.062 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.043 | ||
| RF | 9181 | 0.001 | 0.001 | 0.944 | 0.013 | ||
| RF | 92 | 0.077 | 0.008 | 0.661 | 0.192 | ||
| RF | 184 | 0.015 | 0.001 | 0.927 | 0.142 | ||
| RF | 919 | 0.001 | 0.001 | 0.953 | 0.063 | ||
| RF | 1837 | -0.001 | 0.001 | 0.951 | 0.044 | ||
| RF | 18361 | 0.001 | 0.001 | 0.953 | 0.014 | ||
| Ridge | 23 | 0.137 | 0.023 | 0.446 | 0.247 | ||
| Ridge | 46 | 0.044 | 0.003 | 0.754 | 0.134 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.951 | 0.069 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.945 | 0.049 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.945 | 0.016 | ||
| Ridge | 46 | 0.09 | 0.011 | 0.618 | 0.210 | ||
| Ridge | 92 | 0.021 | 0.002 | 0.916 | 0.159 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.951 | 0.078 | ||
| Ridge | 919 | -0.001 | 0.001 | 0.952 | 0.056 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.953 | 0.017 | ||
| Ridge | 92 | 0.042 | 0.004 | 0.877 | 0.205 | ||
| Ridge | 184 | 0.002 | 0.002 | 0.957 | 0.164 | ||
| Ridge | 919 | -0.002 | 0.001 | 0.950 | 0.075 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.946 | 0.052 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.949 | 0.017 |
| Algorithm | Event rate | N | N event | Bias | MSE | Coverage (ESE) | Width (ESE) |
| GLM | 23 | 0.002 | 0.001 | 0.918 | 0.021 | ||
| GLM | 46 | 0.001 | 0.001 | 0.925 | 0.003 | ||
| GLM | 230 | 0.001 | 0.001 | 0.955 | 0.001 | ||
| GLM | 460 | 0.001 | 0.001 | 0.948 | 0.001 | ||
| GLM | 4591 | 0.001 | 0.001 | 0.896 | 0.000 | ||
| GLM | 46 | 0.001 | 0.001 | 0.926 | 0.006 | ||
| GLM | 92 | 0.001 | 0.001 | 0.939 | 0.003 | ||
| GLM | 460 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| GLM | 919 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| GLM | 9181 | 0.001 | 0.001 | 0.947 | 0.000 | ||
| GLM | 92 | 0.001 | 0.001 | 0.941 | 0.007 | ||
| GLM | 184 | 0.001 | 0.001 | 0.938 | 0.004 | ||
| GLM | 919 | 0.001 | 0.001 | 0.951 | 0.002 | ||
| GLM | 1837 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| GLM | 18361 | 0.001 | 0.001 | 0.929 | 0.000 | ||
| RF | 23 | -0.004 | 0.004 | 0.021 | 0.004 | ||
| RF | 46 | -0.006 | 0.006 | 0.000 | 0.003 | ||
| RF | 230 | -0.005 | 0.005 | 0.000 | 0.001 | ||
| RF | 460 | -0.006 | 0.006 | 0.000 | 0.001 | ||
| RF | 4591 | 0.001 | 0.001 | 0.951 | 0.000 | ||
| RF | 46 | 0.001 | 0.001 | 0.949 | 0.005 | ||
| RF | 92 | 0.001 | 0.001 | 0.952 | 0.003 | ||
| RF | 460 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| RF | 919 | 0.001 | 0.001 | 0.953 | 0.001 | ||
| RF | 9181 | 0.001 | 0.001 | 0.948 | 0.000 | ||
| RF | 92 | 0.001 | 0.001 | 0.954 | 0.007 | ||
| RF | 184 | 0.001 | 0.001 | 0.950 | 0.005 | ||
| RF | 919 | 0.001 | 0.001 | 0.949 | 0.002 | ||
| RF | 1837 | 0.001 | 0.001 | 0.947 | 0.001 | ||
| RF | 18361 | 0.001 | 0.001 | 0.943 | 0.000 | ||
| Ridge | 23 | 0.001 | 0.001 | 0.952 | 0.004 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.951 | 0.003 | ||
| Ridge | 230 | 0.001 | 0.001 | 0.944 | 0.001 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| Ridge | 4591 | 0.001 | 0.001 | 0.930 | 0.000 | ||
| Ridge | 46 | 0.001 | 0.001 | 0.955 | 0.005 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.951 | 0.003 | ||
| Ridge | 460 | 0.001 | 0.001 | 0.950 | 0.001 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.946 | 0.001 | ||
| Ridge | 9181 | 0.001 | 0.001 | 0.947 | 0.000 | ||
| Ridge | 92 | 0.001 | 0.001 | 0.955 | 0.006 | ||
| Ridge | 184 | 0.001 | 0.001 | 0.946 | 0.004 | ||
| Ridge | 919 | 0.001 | 0.001 | 0.952 | 0.002 | ||
| Ridge | 1837 | 0.001 | 0.001 | 0.949 | 0.001 | ||
| Ridge | 18361 | 0.001 | 0.001 | 0.931 | 0.000 |
S1.5 Determining fixed tuning parameters
To determine the values of fixed tuning parameters, we ran a small pilot simulation study. Here, we again varied and considered the three event rates defined in the main manuscript. In this pilot simulation, we bootstrap sampled outcomes along with covariates; we do not expect that this made a meaningful difference in the behavior of AUC. This investigation consisted of two parts: 1) nested cross-validation for tuning parameter selection (inner folds) and estimating AUC (outer folds); and 2) a small number of simulations, with relatively few replications, to assess if there was a large difference between the cross-validated AUC when the selected fixed tuning parameters were used and the cross-validated AUC obtained when we used 20-fold cross-validation to select the tuning parameter in the sampled training set. We referred to the second part of this fixed tuning parameter investigations as pilot simulations. For ridge regression, we used internal tuning from the R function cv.glmnet. For random forests, we set the number of trees to be 250 for , 100 for , and 50 for and tuned over grids of minimum node size given in Table S23. We assessed how variable the selected tuning parameters were across simulation replicates, and determined that the variability was small enough to justify fixing the tuning parameters as described above.
| Training set size | Grid of minimum node sizes |